### v1

# Enhancing Test Time Adaptation with Few-shot Guidance

Abstract
Deep neural networks often encounter significant performance drops while facing with domain shifts between training (source) and test (target) data. To address this issue, Test Time Adaptation (TTA) methods have been proposed to adapt pre-trained source model to handle out-of-distribution streaming target data. Although these methods offer some relief, they lack a reliable mechanism for domain shift correction, which can often be erratic in real-world applications. In response, we develop Few-Shot Test Time Adaptation (FS-TTA), a novel and practical setting that utilizes a few-shot support set on top of TTA. Adhering to the principle of few inputs, big gains, FS-TTA reduces blind exploration in unseen target domains. Furthermore, we propose a two-stage framework to tackle FS-TTA, including (i) fine-tuning the pre-trained source model with few-shot support set, along with using feature diversity augmentation module to avoid overfitting, (ii) implementing test time adaptation based on prototype memory bank guidance to produce high quality pseudo-label for model adaptation. Through extensive experiments on three cross-domain classification benchmarks, we demonstrate the superior performance and reliability of our FS-TTA and framework.
Introduction
In recent years, deep neural networks have exhibited remarkable capabilities in representation learning. However, their performance relies heavily on the assumption that the distributions of training (source) and test (target) data are identical (Long et al. 2015; Ganin and Lempitsky 2015; Li et al. 2017). In real-world deployment, such a distribution shift is inevitable, as it is practically impossible to collect and annotate data for all possible environments in advance of training. Besides, this distribution shift can significantly degrade the performance of the deployed source model.
To address the aforementioned issues, numerous studies have proposed solutions via domain adaptation (Long et al. 2015; Tzeng et al. 2017; Zhang et al. 2019; Xiao and Zhang 2021; Xin et al. 2023) and domain generalization (Volpi et al. 2018; Zhou et al. 2021; Kim et al. 2021; Sicilia, Zhao, and Hwang 2023). While these approaches have demonstrated impressive performance gains on realistic benchmarks, a considerable gap remains between their problem settings and practical application scenarios. Domain adaptation relies on the impractical assumption that target domain data are available and participate in the source training process. In contrast, domain generalization aims to directly enhance the generalization of the source model without exploring the target domain data, even if they can be obtained during the test time.
Figure 1: Test Time Adaptation (TTA) vs. Few-Shot Test Time Adaptation (FS-TTA). FS-TTA utilizes labeled few-shot target data in addition to TTA. The results for TTA are derived from the performance of TENT (Wang et al. 2021) on the OfficeHome (Venkateswara et al. 2017).
In order to overcome these limitations of domain adaptation/generalization and protect the privacy of the source data, TENT (Wang et al. 2021) introduces fully test time adaptation (TTA). TTA aims to adapt a pre-trained source model to the target domain using input mini-batch data during the test time, without relying on source data or supervision. TTA is particularly focused on an online setting, where the model must adapt and make predictions immediately upon receiving each batch of potentially non-independent and identically distributed (non-i.i.d.) target samples. To serve this purpose, TENT employs test-time entropy minimization to reduce the generalization error on shifted target data. Additionally, extensive research has sought to improve TTA through various approaches such as pseudo-labeling (Iwasawa and Matsuo 2021; Wang, Zhang et al. 2023), consistency regularization (Boudiaf et al. 2022), and anti-forgetting regularization (Niu et al. 2022). While these methods can perform model adaptation during the test time, they encounter three primary challenges:
1) Domain shift correction: The certainty of TTA methods in addressing domain shifts effectively without utilizing target labels is questionable. The t-SNE visualization in Fig. 5 clearly illustrates this point, where we observe that the feature distribution exhibits negligible change following the adaptation process with TENT. This suggests that TTA methods may struggle to effectively adjust to new domain characteristics in the complete absence of target labels, which could provide essential guidance for adaptation.
2) Generalizability: The effectiveness of TTA methods varies across different scenarios. In some cases, they might even underperform compared to the pre-trained source model without any adaptation, as illustrated in Fig. 1 (Source Model vs. TENT). This variability indicates that the generalization performance of TTA methods is not particularly strong and can be influenced by various factors, including the domain shift and the specific characteristics of the model and dataset involved.
3) Data reliance: The success of TTA methods heavily relies on the availability and quality of unlabeled mini-batch data from the target domain. This reliance presents a challenge, as the adaptation process is directly influenced by the representativeness, quantity, and quality of the available unlabeled data. In scenarios where high-quality, relevant unlabeled data is scarce or not fully representative of the entire target domain, TTA methods may face difficulties in achieving optimal performance, highlighting a major limitation in their application across various real-world settings.
Figure 2: The performance impact of one-shot.
The fundamental reason for these challenges is the blind exploration of the target domain without any supervision information when domain shift exists. Actually, in practical scenarios, it’s quite feasible to obtain few-shot labeled samples from the target domain. Consequently, this leads us to ask: If given little target domain supervisory information, could the adaptation performance be improved?
To answer this question, we test the one-shot situation, as shown in Fig. 2. Specifically, we use one sample per class to fine-tune the source model with cross-entropy loss. We find that the performance is easily improved compared to TENT, which shows that little supervision information can be more effective than a large amount of unsupervised information.
Based on the aforementioned findings, we introduce Few-Shot Test Time Adaptation (FS-TTA), encapsulating the concept of few inputs, big gains. As illustrated in Fig. 1, by integrating a few-shot support set from the target domain prior to model adaptation, FS-TTA effectively reduces domain shift while retaining the source-free and online characteristics inherent to TTA. It is noted that such setting is particularly beneficial in scenarios where precision and reliability are paramount, such as medical image analysis and autonomous driving, even spending a few extra labeling costs.
To solve the FS-TTA, we develop an effective framework. For domain shift correction, we first fine-tune the pre-trained source model with the few-shot support set, fostering initial adaptation to the target. To prevent overfitting, we propose Feature Diversity Augmentation (FDA) to generate new features. During the test time, we employ a self-training strategy, which involves assigning pseudo-labels to unlabeled online mini-batches and using these labels to further update the model online. Furthermore, in order to reduce the impact of noisy pseudo-labels on the model, we propose Entropy Filter and Consistency Filter. The former filters out high-entropy samples with low confidence, and the latter is achieved through dual-branch prediction consistency. The experimental results across various cross-domain image recognition datasets show that our FS-TTA method significantly surpasses the performance of state-of-the-art methods and other baselines. To sum up, our main contributions are as follows:
  • New research direction: We highlight the setting of Few-Shot Test Time Adaptation (FS-TTA), in which an additional few-shot support set is available prior to test time adaptation. Thanks to the few-shot samples, domain shift can be reduced more effectively.
  • Novel framework: We propose a carefully crafted framework to tackle FS-TTA, including fine-tuning the pre-trained source model with feature diversity augmentation module and performing test time adaptation via high-quality pseudo-labeled samples in self-training manner.
  • Superior performance: We provide empirical results on various cross-domain classification benchmarks to demonstrate the superiority of our framework. In comparison to the state-of-the-art method in TTA, our method achieves improvements of 2.0% on PACS, 7.8% on OfficeHome, and 3.9% on DomainNet.
Related Work
Test Time Adaptation
Test Time Adaptation (TTA) strives to adapt the pre-trained source model during the test time to alleviate distribution shifts. Previous TTA methods address distribution alignment by incorporating self-supervised tasks (Sun et al. 2020). Although effective, TTA necessitates access to training samples and modifies the training process. To surmount this challenge, TENT (Wang et al. 2021) introduces the concept of fully test-time adaptation, utilizing only target data. TENT adapts the batch normalization layer by minimizing the entropy of model predictions. Distinct from TENT, recent works (Schneider et al. 2020; Nado et al. 2020) estimate the mean and variance of the activation in each batch normalization on the forthcoming test mini-batch. Consequently, LAME (Boudiaf et al. 2022) and EATA (Niu et al. 2022) concentrate on tackling the issue of catastrophic forgetting as the source model undergoes continuous updates. TSD (Wang, Zhang et al. 2023) integrates self-training into TTA, selecting high-quality test samples to fit the model.
Table 1: Comparison with various adaptation settings, where $s$ and $t$ denote source domain and target domain, respectively. $L^{d}$ and $U^{d}$ denote labeled datasets and unlabeled datasets from domain $d$. “Online” means that adaptation can predict a batch of incoming test samples immediately. “$k$” represents the number of samples per class. “C” indicates the number of classes for the target domain.
Setting Source-free Training inputs Online
Source domain(s) Target domain Size of available target data
[…]
Comparisons with Other Settings
We compare Few-Shot Test Time Adaptation (FS-TTA) with similar problem settings (details are in the appendix), as illustrated in Tab. 1.
  • Compared with Domain Generalization, FS-TTA eliminates the need for access to source data, thereby safeguarding the privacy of the source data. Moreover, FS-TTA facilitates adaptation to the downstream target domain via model parameter updates.
  • Compared with Source-Free Domain Adaptation, FS-TTA obviates the need to acquire all target domain data simultaneously and can dynamically update the model online, contingent upon incoming mini-batches.
  • Compared with Few-Shot Transfer Learning, FS-TTA can not only utilize a small number of target samples, but can also further update the model using online mini-batch target data during the test time.
  • Compared with Test Time Adaptation, FS-TTA capitalizes on an auxiliary small set of samples from the target, allowing the pre-trained source model to adapt more swiftly and proficiently to the target domain. Furthermore, FS-TTA demonstrates exceptional performance in managing circumstances involving substantial domain shifts.
Method
Problem Setting
Considering a typical scenario where a source model $f_{\theta_{s}}$, is equipped with parameters $\theta_{s}$ and trained on source datasets $\mathcal{D}_{s_{1}},\mathcal{D}_{s_{2}},...,\mathcal{D}_{s_{n}}$, our goal is to adapt this pre-trained source model to a target domain $D_{t}$ without accessing source data. We postulate the existence of a small, labeled support set $S=\{(s_{i},y_{i})\}$ derived from $D_{t}$, where $s_{i}$ represents the image, and $y_{i}$ is its corresponding label. During the test time, mini-batches unlabeled target samples online arrive. Few-Shot Test Time Adaptation (FS-TTA) strives to effectively adapt the pre-trained source model $f_{\theta_{s}}$, utilizing the labeled support set $S$ and unlabeled online mini-batch to tackle the domain shift challenge. Notably, the support set $S$ could be offline obtained before the test time.
Stage I: Model Adaptation via Fine-Tuning
To significantly and swiftly enhance the initialization performance of the pre-trained source model in the target domain and minimize domain shifts, we design to fine-tune the pre-trained source model using the few-shot support set. Given the limited number of samples per class, there is a potential risk of overfitting during the fine-tuning process. To mitigate this, we introduce the Feature Diversity Augmentation (FDA) module, which generates new features by mixing statistics. Ultimately, we use a supervised classification loss to fine-tune the pre-trained source model. This entire procedure is illustrated in Stage I of Fig. 3.
Figure 3: Illustration of our two-stage framework. In Stage I, we employ the few-shot support set to fine-tune the source model. To prevent overfitting, we propose FDA module. In Stage II, we maintain a prototype memory bank to guide test time adaptation. In order to update the prototype memory bank and model with effective samples, we propose the entropy filter and consistency selection modules.
Feature Diversity Augmentation.
Prior research (Zhou et al. 2021) has demonstrated a significant association between feature statistics and image style, which is intricately linked to data distribution within the field of computer vision. Inspired by their discoveries, we propose Feature Diversity Augmentation (FDA), which enhances feature diversity and mitigates the risk of overfitting during fine-tuning.
FDA is incorporated between layers (blocks) in the pre-trained source backbone, as depicted in Fig. 3. More specifically, FDA mixes the feature statistics of two random samples to generate new features. The computations within FDA module can be summarized in three steps. Firstly, given two feature maps $f_{i}$ and $f_{j}$ from the support set, we compute their feature statistics ($\mu_{i},\sigma_{i}$) and ($\mu_{j},\sigma_{j}$). Secondly, FDA generates the mixtures of feature statistics:
$\boldsymbol{\gamma}_{\text{mix }}=\lambda\sigma_{i}+(1-\lambda)\sigma_{j},$ (1)
$\boldsymbol{\beta}_{\text{mix }}=\lambda\mu_{i}+(1-\lambda)\mu_{j}.$ (2)
In this case, $\lambda$ denotes the mixing ratio coefficient. Ultimately, the mixtures of feature statistics are applied to the feature map $f_{i}$ via instance normalization:
$f_{i}^{\prime}=\boldsymbol{\gamma}_{\operatorname{mix}}\odot\frac{f_{i}-\mu_{i}}{\sigma_{i}}+\boldsymbol{\beta}_{\operatorname{mix}},$ (3)
where $f_{i}^{\prime}$ represents the newly generated feature map.
Fine-Tuning Source Model.
To enhance the adaptation of the pre-trained source model to the target, we employ the few-shot support set to fine-tune the model with the FDA module. Specifically, the few-shot support set is processed through $f_{\theta_{s}}$ to minimize a supervised loss, defined as:
$\mathcal{L}_{\text{cls}}=-\sum_{i=1}^{k*C}\mathcal{H}\left(y_{i},p\left(\hat{y_{i}}\mid\boldsymbol{s}_{i}\right)\right),$ (4)
where $\mathcal{H}(\cdot)$ is the cross-entropy loss. The term $y_{i}$ is the ground-truth label of $s_{i}$, indicating one of sample from few-shot support set, and $C$ represents categories of the target.
Table 2: Compared with test time adaptation methods on three datasets with ResNet50 backbone.
OfficeHome PACS DomainNet
Method Art Clip Prod Real Avg. Art Cart Phot Sket Avg. Avg.
Test time adaptation methods
ERM (Vapnik 1999) 60.7 55.7 76.2 76.8 67.4 82.5 80.8 94.0 80.9 84.5 45.2
BN (Nado et al. 2020) 58.2 55.6 75.1 75.5 66.1 83.2 84.9 94.0 77.9 85.0 43.3
TENT (Wang et al. 2021) 60.6 58.7 76.5 76.1 68.0 85.2 86.7 94.9 82.9 87.4 44.7
T3A (Iwasawa and Matsuo 2021) 61.2 56.7 78.0 77.3 68.3 84.0 82.3 95.0 82.7 86.0 46.1
ETA (Niu et al. 2022) 58.4 55.8 75.2 75.5 66.2 83.2 84.9 94.0 77.9 85.0 46.1
LAME (Boudiaf et al. 2022) 58.7 55.6 75.1 75.4 66.2 84.9 85.5 95.0 80.9 86.6 43.2
TSD (Wang, Zhang et al. 2023) 62.3 57.5 77.5 77.5 68.7 87.6 88.7 96.1 85.0 89.4 47.7
Fine-tuning + Test time adaptation methods
FT+TENT (Wang et al. 2021) 68.8 65.5 79.8 78.5 73.2 87.0 86.9 95.2 83.6 88.2 45.4
FT+TSD (Wang, Zhang et al. 2023) 70.5 65.1 80.3 79.2 73.8 88.3 88.6 96.5 85.9 89.8 48.5
FS-TTA 73.2 68.3 83.0 81.6 76.5 90.4 89.7 97.6 87.8 91.4 51.6
$\Delta_{up}$ over TSD (+10.9$\uparrow$ (+10.8$\uparrow$ (+5.5$\uparrow$ (+4.1$\uparrow$ (+7.8$\uparrow$ (+2.8$\uparrow$ (+1.0$\uparrow$ (+1.5$\uparrow$ (+2.8$\uparrow$ (+2.0$\uparrow$ (+3.9$\uparrow$
Stage II: Test Time Adaptation
During this stage, a mini-batch of unlabeled samples, denoted as $x=\{x_{1},x_{2},..,x_{B}\}$, online arrives. The central concept of this stage is to employ a self-training strategy to update the fine-tuned source model online, enabling it to fully adapt to the target domain. This involves assigning pseudo-labels to unlabeled online mini-batches and using these labels to further update the model. Thus, we first generate the pseudo-labels by $\hat{y_{i}}=\text{argmax}(p_{i})$ for $x_{i}$, where $p_{i}$ is the prediction logits. However, it is inevitable that there are always some noisy samples are misclassified, leading to wrong pseudo-labels. To address this issue, we propose two modules to produce high quality pseudo-labels. The first is entropy filter, which screens out unreliable samples using Shannon entropy (Shannon 2001). Typically, samples with higher entropy are considered to have lower prediction confidence. The second module is a prototype memory bank classification, which works in tandem with the classifier. The prototype memory bank is used to generate pseudo-labels outside the classifier, according to the nearest class prototype in the feature space. After that, pseudo-labels with consistency prediction is preserved for model adaptation. The entire process is outlined in Stage II of Fig. 3.
Entropy Filter.
To dynamically update the model using online mini-batch target, it is crucial to filter out noisy samples, as they may be assigned to incorrect classes, resulting in inaccurate prototype computation. In this regard, we propose the Entropy Filter, which employs Shannon entropy (Shannon 2001) to select confident samples in the mini-batch. For an sample $x_{i}$, its entropy can be computed as:
$H(p_{i})=-\sum(p_{i})\cdot log(p_{i}).$ (5)
Based on the insights from previous work (Wang et al. 2021), high entropy samples should be filtered out, as lower entropy typically indicates higher accuracy. Consequently, we sort the entropy of all samples in the mini-batch and select the top $\alpha\%$ samples with lower entropy, donated as $\hat{x}=\{\hat{x}_{1},\hat{x}_{2},...,\hat{x}_{|\alpha\%*B|}\}$.
Prototype Memory Bank.
We maintain a prototype memory bank $M=\{m_{1},m_{2},...,m_{C}\}$ to store class prototypes, where $C$ represents categories of the target. The prototype memory bank is initialized with the few-shot support set $S$, defined as:
$m_{c_{0}}=\frac{\sum_{i=1}^{|S|}f_{i}\cdot\mathbbm{1}[y_{i}=c]}{\sum_{i=1}^{|S|}\mathbbm{1}[y_{i}=c]},$ (6)
where $\mathbbm{1}[\cdot]$represents an indicator function, yielding a value of 1 if the argument is true or 0 otherwise, and $m_{c_{0}}$ denotes the initial moment of the $c$-th class prototype. Thanks to the few-shot support, precise guidance can be provided during the initial phase, thereby reducing reliance on the quality of online mini-batch data.
Throughout the test time adaptation process, we persistently update the prototype memory bank by incorporating selected reliable samples with pseudo labels:
$m_{c_{t}}=\beta*m_{c_{t-1}}+(1-\beta)*\frac{\sum_{j=1}^{|\hat{x}|}f_{j}\cdot\mathbbm{1}[\hat{y}_{j}=c]}{\sum_{j=1}^{|\hat{x}|}\mathbbm{1}[\hat{y}_{j}=c]},$ (7)
where $m_{c_{t}}$ represents the $c$-th class prototype at time $t$, and $\beta$ represents the sliding update coefficient.
Test Time Adaptation.
During the test time adaptation, we adopt high-quality pseudo-labeled samples to guide the model update. First, we define the prototype-based classification output as the softmax over the feature similarity to prototypes for class $c$:
$\hat{p}_{j}^{c}=\frac{\exp\left(\operatorname{sim}\left(f_{j},m_{c}\right)\right)}{\sum_{c=1}^{C}\exp\left(\operatorname{sim}\left(f_{j},m_{c}\right)\right)},$ (8)
where $sim(\cdot,\cdot)$ represents cosine similarity. Subsequently, we propose that, for a reliable sample, the outputs of the fine-tuned model and prototype-based classification should be similar. Therefore, we propose the consistency filter to identify incorrect predictions. This strategy can be implemented through a filter mask for samples $x_{j}$ as follows:
$\mathcal{M}_{j}=\mathbbm{1}\left[\arg\max p_{j}=\arg\max\hat{p}_{j}\right].$ (9)
Ultimately, we can update the model using reliable samples, and the loss can be formulated as follows:
$\mathcal{L}_{online}=\frac{{\textstyle\sum_{j=1}^{|\hat{x}|}}\mathcal{H}_{j}*\mathcal{M}_{j}}{{\textstyle\sum_{j=1}^{|\hat{x}|}}\mathcal{M}_{j}}.$ (10)
It’s noteworthy that our self-training process does not involve specifying any threshold, which enhances the model’s generalizability.
Experiment
Experimental Settings
Dataset.
To test the effectiveness of our setting and method, we experiment on three cross-domain benchmarks. PACS (Li et al. 2017) consists of four distinct domains: Art, Cartoon, Photo, and Sketch, and comprises a total of 9,991 images and 7 classes. Office-Home (Venkateswara et al. 2017) consists of 15,588 images from four distinct image domains, namely Real World, Clipart, Art, and Product, with 65 classes. DomainNet (Peng et al. 2019) is a large-scale dataset that comprises six domains, denoted as $d\in${Clipart, Infograph, Painting, Quickdraw, Real, Sketch}, with a total of 586,575 images and 345 classes.
Implementation Details.
In the main experiments, we utilize ResNet-50 (He et al. 2016) pre-trained on ImageNet-1k (Russakovsky et al. 2015) as our backbone, which is widely used in test time adaptation literature. For source model training, we adhere to the leave-one-domain-out protocol, as recommended by previous work  (Wang, Zhang et al. 2023; Zhou et al. 2021). We employ the Adam optimizer with a learning rate of $5e^{-5}$. For few-shot test time adaptation, we also employ the Adam optimizer and set the batch size to The few-shot support set typically selects 5 to 16 samples per class, depending on the difficulty of the target. We carry out all experiments on NVIDIA V100 GPUs.
Baselines.
We compare our method with test time adaptation methods, BN (Nado et al. 2020), TENT (Wang et al. 2021), ETA (Niu et al. 2022), T3A (Iwasawa and Matsuo 2021), LAME (Boudiaf et al. 2022), and TSD (Wang, Zhang et al. 2023). Additionally, we create new baselines by combining fine-tuning with TTA methods for a more comprehensive comparison. We also compare with some methods in domain generalization and source-free domain adaptation, including DNA (Chu et al. 2022), PCL (Yao et al. 2022), SWAD (Cha et al. 2021), and F-mix (Kundu et al. 2022).
Table 3: Compared with existing DG and SFDA methods on OfficeHome and DomainNet.
OfficeHome DomainNet
Method Art Clip Prod Real Avg. Clip Info Pain Quic Real Sket Avg.
Domain generalization methods
ERM (Vapnik 1999) 60.7 55.7 76.2 76.8 67.4 64.8 22.1 51.8 13.8 64.7 54.0 45.2
DNA (Chu et al. 2022) 67.7 57.7 78.9 80.5 71.2 66.1 23.0 54.6 16.7 65.8 56.8 47.2
PCL (Yao et al. 2022) 67.3 59.9 78.7 80.7 71.6 67.9 24.3 55.3 15.7 66.6 56.4 47.7
SWAD (Cha et al. 2021) 66.1 57.7 78.4 80.2 70.6 66.1 22.4 53.6 16.3 65.5 56.2 46.7
Source-free domain adaptation methods
F-mix (Kundu et al. 2022) 72.6 67.4 85.9 83.6 77.4 75.4 24.6 57.8 23.6 65.8 58.5 51.0
FS-TTA 73.2 68.3 83.0 81.6 76.5 68.6 30.8 56.4 24.2 69.1 60.2 51.6
SWAD + FS-TTA 77.4 71.1 86.4 84.2 79.8 - - - - - - -
Performance Comparisons
Comparison with TTA methods.
Tab. 2 details the comparison results between our method and various TTA methods on the Office-Home and PACS datasets, as well as the final results of DomainNet (detailed in Tab. 3). We observe that our method achieves state-of-the-art performance.
Primarily, our approach exhibits a significant enhancement in performance compared to the source model (ERM). Our FS-TTA achieves improvements across all four tasks on Office-Home, with gains of 12.5% (Art), 12.6% (Clipart), 6.8% (Product), and 4.8% (Real), respectively. Notably, our method demonstrates more substantial improvement on the more challenging tasks (e.g., Art and Clipart), confirming that FS-TTA is more friendly for large domain shifts. On the other two datasets, we observe average performance increments of 6.9% (PACS) and 6.4% (DomainNet).
Moreover, our method outperforms the state-of-the-art TTA method, TSD, with average performance increments of 2.0% (PACS), 7.8% (Office-Home), and 3.9% (DomainNet). The lesser improvement in PACS can be attributed to its lower complexity, while our method shows superior performance on the more challenging Office-Home and DomainNet datasets. This significant improvement benefits from our effective utilization of few-shot target information, including the FDA module and initializing the prototype memory bank. The performance of some TTA methods, such as ETA and LAME, does not meet the expected standards on Office-Home and other datasets. In fact, they even exhibit inferior performance compared to the source model on certain tasks (e.g., Art, Product, and Real), which highlights the limitations of TTA and the necessity of few-shot target samples. In conclusion, our FS-TTA demonstrates a notable advantage in tasks that closely resemble real-world scenarios and provides a significant boost in performance with minimal additional computational overhead.
Finally, for a more comprehensive comparison with TTA methods, we construct new baselines, which combine fine-tuning with TTA methods (e.g., TENT and TSD). According to the results in Tab. 2, our method shows an average (three datasets) improvement of 4.2% over Fine-Tuning+TENT, and 2.4% over Fine-Tuning+TSD, which demonstrates the superiority of our method over existing TTA techniques in migrating to few-shot TTA scenarios. This result benefits from our unique FDA module design in Fine-Tuning and the use of few-shot support set to initialize prototype memory bank.
Refer to captionRefer to caption(a) Different stage(b) Different $\alpha$
Figure 4: (a) effectiveness analysis about two-stage framework and (b) sensitivity analysis about parameter $\alpha$.
Comparison with DG/SFDA methods.
The above experiments mainly focus on TTA, which aims to adapt the model during the test time. A natural question arises: How about our method compared with domain generalization (DG) or source-free domain adaptation (SFDA) methods?
To answer this question, we compare our method with several methods in DG and SFDA. The results of Office-Home dataset are shown in Tab. 3. It can be seen that our method outperforms the state-of-the-art methods in DG, such as SWAD and PCL. Furthermore, equipped with SWAD (SWAD+FS-TTA), our method achieves 79.8% accuracy. This result benefits from our adaptation of the model during the test time. In comparison to advanced SFDA methods, FS-TTA still achieves satisfactory results. It is worth noting that FS-TTA is more flexible in real-world scenarios than SFDA since it adapts the target data in an offline manner, requiring more training loops and resources. The results of DomainNet are shown in Tab. 3. The overall performance of FS-TTA outperforms the SFDA methods, suggesting that FS-TTA is more adept at handling challenging tasks.
Ablation Study
Figure 5: The t-SNE feature visualization of (a) ERM, (b) TENT, (c) LAME, and (d) our method.
Effectiveness of two-stage framework.
Our proposed method consists of two stages, with the individual contributions of each stage presented in Fig. 4(a). Compared to the baseline source model, Stage I of our approach achieves an average improvement of 6.6% on the Office-Home. This highlights the effectiveness of our fine-tuning strategy, which employs a mixture of statistics between samples, validating its suitability for the target domain. Our test time adaptation method, which relies on class prototype memory bank guidance during Stage II, adds an extra 2.5% performance enhancement. As a result, our two-stage framework establishes itself as a robust foundation for the Few-Shot Test Time Adaptation setting, demonstrating its considerable potential in enabling online model adaptation in real-world situations where labeled data is scarce.
Sensitivity to $\alpha$.
The parameter $\alpha$ represents the proportion of each batch that is selected through an entropy filter to update the prototype memory bank and the model. To evaluate the impact of $\alpha$, we conduct an experimental analysis on the Office-Home dataset by assigning $\alpha$ to 0, 0.3, 0.6, and 1, respectively. The results, as shown in Fig. 4(b), demonstrate that $\alpha>0$ yields performance improvements compared to $\alpha=0$ (the source model), highlighting the effectiveness of our proposed framework. Furthermore, $\alpha=0.3$ and $\alpha=0.6$ perform better than $\alpha=1$ (no filter), indicating the effectiveness of our entropy filter strategy.
Qualitative analysis by t-SNE visualization.
We present t-SNE visualizations to compare the feature representations of the pre-trained source model (ERM), test time adaptation methods (TENT and LAME), and our proposed method, as illustrated in Fig. 5. The learned features of the pre-trained source model on the target domain are not well-separated due to the significant domain gap, as shown in Fig. 5(a). Additionally, we can observe no considerable feature distribution changes on the target domain after adaptation with TENT and LAME methods, as shown in Fig. 5(b) and Fig. 5(c). In contrast, our method produces more uniform and aligned feature distribution after adapting to the target domain, as shown in Fig. 5(d).
Efficiency analysis.
In our main experiments, we opt for a mini-batch size of 64. To examine the variations in performance and computational efficiency with different batch size during test-time adaptation, we conduct a series of analytical experiments. As shown in Fig. 6(a), we observe that accuracy experiences a gradual increase as the batch size incrementally grows, reaching a plateau around a batch size of 64. In contrast, as shown in Fig. 6(b), running time exhibits a decreasing trend as the batch size grows. However, beyond a batch size of 64, the running time appears to stabilize. Consequently, for real-world applications aiming to achieve a trade-off between accuracy and computational efficiency, we suggest a batch size in the vicinity of 64.
Refer to captionRefer to caption(a) Accuracy (batch size)(b) Running Time
Figure 6: During the test time adaptation, we analyze:(a) accuracy and (b) running time, across various batchsize.
Figure 7: Ablation experiments on shot size.
Ablation experiments on shot size.
To elucidate the impact of the number of k-shots on our method, we carry out additional ablation experiments within the Office-Home dataset. The findings, illustrated in Figure 7, indicate a significant performance enhancement when the shot size ranges from 1 to 10, demonstrating a rapid performance ascension in this few-shot regime. Remarkably, even minimal shot sizes such as 1-shot and 3-shot exhibit substantial effectiveness. For instance, the 3-shot configuration achieves a 3.8% performance improvement over the TSD.
Conclusion
In this work, we introduce Few-Shot Test Time Adaptation (FS-TTA), a novel setting that diverges from traditional TTA by leveraging the few-shot support set to improve adaptation to the target. To tackle FS-TTA, we propose an effective framework, which involves employing the few-shot support set to fine-tune the pre-trained source model and maintaining a prototype memory bank to guide the test time adaptation. Results on three cross-domain benchmarks demonstrate the superior performance and reliability of our method. Looking ahead, we aspire to expand FS-TTA beyond current scope by investigating potential real-world tasks, instead of limiting to image recognition.

### v2

@@
-Enhancing Test Time Adaptation with Few-shot Guidance
+
 Abstract
-Deep neural networks often encounter significant performance drops while facing with domain shifts between training (source) and test (target) data.
-To address this issue, Test Time Adaptation (TTA) methods have been proposed to adapt pre-trained source model to handle out-of-distribution streaming target data.
-Although these methods offer some relief, they lack a reliable mechanism for domain shift correction, which can often be erratic in real-world applications.
-In response, we develop Few-Shot Test Time Adaptation (FS-TTA), a novel and practical setting that utilizes a few-shot support set on top of TTA.
-Adhering to the principle of few inputs, big gains, FS-TTA reduces blind exploration in unseen target domains.
-Furthermore, we propose a two-stage framework to tackle FS-TTA, including (i) fine-tuning the pre-trained source model with few-shot support set, along with using feature diversity augmentation module to avoid overfitting, (ii) implementing test time adaptation based on prototype memory bank guidance to produce high quality pseudo-label for model adaptation.
-Through extensive experiments on three cross-domain classification benchmarks, we demonstrate the superior performance and reliability of our FS-TTA and framework.
-Introduction
+Deep neural networks often experience significant performance degradation when encountering domain shifts between training (source) and test (target) data.
+To mitigate this issue, Test Time Adaptation (TTA) methods have been introduced to adapt a pre-trained source model for handling out-of-distribution streaming target data.
+While these methods provide some improvement, they often lack a dependable mechanism for domain shift correction, which can be unpredictable in real-world applications.
+To address this challenge, we propose the Few-Shot Test Time Adaptation (FS-TTA), a novel and practical setting that builds upon TTA by leveraging a few-shot support set.
+Following the principle of few inputs, big gains, FS-TTA effectively reduces the uncertainty of blind adaptation in previously unseen target domains.
+Furthermore, we introduce a two-stage framework to effectively tackle FS-TTA.
+The first stage involves fine-tuning the pre-trained source model using the few-shot support set while incorporating a feature diversity augmentation module to prevent overfitting.
+The second stage utilizes test time adaptation guided by a prototype memory bank, which facilitates the generation of high-quality pseudo-labels for model adaptation.
+Extensive experiments conducted on three cross-domain classification benchmarks validate the effectiveness, reliability, and superior performance of our proposed FS-TTA and framework.
+keywords:
+Test Time Adaptation, Few-Shot, Domain Shift, Transfer Learning.
+Figure 1: Test Time Adaptation (TTA) vs.
+Few-Shot Test Time Adaptation (FS-TTA).
+FS-TTA incorporates a small number of labeled target samples, which can be easily collected offline before deployment with minimal annotation effort, in addition to the unlabeled target data used in TTA.
+The results for TTA are based on the performance of TENT (Wang et al, 2021) on the OfficeHome (Venkateswara et al, 2017).
+1 Introduction
 In recent years, deep neural networks have exhibited remarkable capabilities in representation learning.
-However, their performance relies heavily on the assumption that the distributions of training (source) and test (target) data are identical (Long et al.
-2015; Ganin and Lempitsky 2015; Li et al.
-2017).
+However, their performance relies heavily on the assumption that the distributions of training (source) and test (target) data are identical (Long et al, 2015; Ganin and Lempitsky, 2015; Li et al, 2017).
 In real-world deployment, such a distribution shift is inevitable, as it is practically impossible to collect and annotate data for all possible environments in advance of training.
 Besides, this distribution shift can significantly degrade the performance of the deployed source model.
-To address the aforementioned issues, numerous studies have proposed solutions via domain adaptation (Long et al.
-2015; Tzeng et al.
-2017; Zhang et al.
-2019; Xiao and Zhang 2021; Xin et al.
-2023) and domain generalization (Volpi et al.
-2018; Zhou et al.
-2021; Kim et al.
-2021; Sicilia, Zhao, and Hwang 2023).
+To address the aforementioned issues, numerous studies have proposed solutions via domain adaptation (Long et al, 2015; Tzeng et al, 2017; Zhang et al, 2019; Xiao and Zhang, 2021; Xin et al, 2023) and domain generalization (Volpi et al, 2018; Zhou et al, 2021; Kim et al, 2021; Sicilia et al, 2023).
 While these approaches have demonstrated impressive performance gains on realistic benchmarks, a considerable gap remains between their problem settings and practical application scenarios.
@@
 In contrast, domain generalization aims to directly enhance the generalization of the source model without exploring the target domain data, even if they can be obtained during the test time.
-Figure 1: Test Time Adaptation (TTA) vs.
-Few-Shot Test Time Adaptation (FS-TTA).
-FS-TTA utilizes labeled few-shot target data in addition to TTA.
-The results for TTA are derived from the performance of TENT (Wang et al.
-2021) on the OfficeHome (Venkateswara et al.
-2017).
-In order to overcome these limitations of domain adaptation/generalization and protect the privacy of the source data, TENT (Wang et al.
-2021) introduces fully test time adaptation (TTA).
+In order to overcome these limitations of domain adaptation/generalization and protect the privacy of the source data, TENT (Wang et al, 2021) introduces fully test time adaptation (TTA).
 TTA aims to adapt a pre-trained source model to the target domain using input mini-batch data during the test time, without relying on source data or supervision.
@@
 To serve this purpose, TENT employs test-time entropy minimization to reduce the generalization error on shifted target data.
-Additionally, extensive research has sought to improve TTA through various approaches such as pseudo-labeling (Iwasawa and Matsuo 2021; Wang, Zhang et al.
-2023), consistency regularization (Boudiaf et al.
-2022), and anti-forgetting regularization (Niu et al.
-2022).
+Additionally, extensive research has sought to improve TTA through various approaches such as pseudo-labeling (Iwasawa and Matsuo, 2021; Wang et al, 2023), consistency regularization (Boudiaf et al, 2022), and anti-forgetting regularization (Niu et al, 2022).
 While these methods can perform model adaptation during the test time, they encounter three primary challenges:
 1) Domain shift correction: The certainty of TTA methods in addressing domain shifts effectively without utilizing target labels is questionable.
-The t-SNE visualization in Fig.
-5 clearly illustrates this point, where we observe that the feature distribution exhibits negligible change following the adaptation process with TENT.
+The t-SNE visualization in Figure 7 clearly illustrates this point, where we observe that the feature distribution exhibits negligible change following the adaptation process with TENT.
 This suggests that TTA methods may struggle to effectively adjust to new domain characteristics in the complete absence of target labels, which could provide essential guidance for adaptation.
 2) Generalizability: The effectiveness of TTA methods varies across different scenarios.
-In some cases, they might even underperform compared to the pre-trained source model without any adaptation, as illustrated in Fig.
-1 (Source Model vs.
+In some cases, they might even underperform compared to the pre-trained source model without any adaptation, as illustrated in Figure 1 (Source Model vs.
 TENT).
@@
 In scenarios where high-quality, relevant unlabeled data is scarce or not fully representative of the entire target domain, TTA methods may face difficulties in achieving optimal performance, highlighting a major limitation in their application across various real-world settings.
-Figure 2: The performance impact of one-shot.
-The fundamental reason for these challenges is the blind exploration of the target domain without any supervision information when domain shift exists.
-Actually, in practical scenarios, it’s quite feasible to obtain few-shot labeled samples from the target domain.
-Consequently, this leads us to ask: If given little target domain supervisory information, could the adaptation performance be improved?
-To answer this question, we test the one-shot situation, as shown in Fig.
-2.
+Figure 2: Performance comparison across different adaptation strategies on OfficeHome.
+One-shot fine-tuning with a single labeled sample per class already surpasses TENT, showing the effectiveness of minimal supervision.
+The fundamental reason for these challenges lies in the blind exploration of the target domain without any supervisory signal under domain shift.
+In practical scenarios, however, it is often feasible to obtain a small number of labeled target samples with minimal annotation cost.
+For instance, a few representative examples can be labeled offline before deployment by domain experts or system users.
+This leads us to ask: If given limited supervisory information from the target domain, could the adaptation performance be improved?
+To answer this question, we test the one-shot situation, as shown in Figure 2.
 Specifically, we use one sample per class to fine-tune the source model with cross-entropy loss.
 We find that the performance is easily improved compared to TENT, which shows that little supervision information can be more effective than a large amount of unsupervised information.
-Based on the aforementioned findings, we introduce Few-Shot Test Time Adaptation (FS-TTA), encapsulating the concept of few inputs, big gains.
-As illustrated in Fig.
-1, by integrating a few-shot support set from the target domain prior to model adaptation, FS-TTA effectively reduces domain shift while retaining the source-free and online characteristics inherent to TTA.
+Based on the aforementioned findings, we introduce the Few-Shot Test Time Adaptation (FS-TTA), encapsulating the concept of few inputs, big gains.
+As illustrated in Figure 1, by integrating a few-shot support set from the target domain prior to model adaptation, FS-TTA effectively reduces domain shift while retaining the source-free and online characteristics inherent to TTA.
 It is noted that such setting is particularly beneficial in scenarios where precision and reliability are paramount, such as medical image analysis and autonomous driving, even spending a few extra labeling costs.
@@
 To sum up, our main contributions are as follows:
-  • New research direction: We highlight the setting of Few-Shot Test Time Adaptation (FS-TTA), in which an additional few-shot support set is available prior to test time adaptation.
-Thanks to the few-shot samples, domain shift can be reduced more effectively.
-  • Novel framework: We propose a carefully crafted framework to tackle FS-TTA, including fine-tuning the pre-trained source model with feature diversity augmentation module and performing test time adaptation via high-quality pseudo-labeled samples in self-training manner.
-  • Superior performance: We provide empirical results on various cross-domain classification benchmarks to demonstrate the superiority of our framework.
-In comparison to the state-of-the-art method in TTA, our method achieves improvements of 2.0% on PACS, 7.8% on OfficeHome, and 3.9% on DomainNet.
-Related Work
-Test Time Adaptation
-Test Time Adaptation (TTA) strives to adapt the pre-trained source model during the test time to alleviate distribution shifts.
-Previous TTA methods address distribution alignment by incorporating self-supervised tasks (Sun et al.
-2020).
-Although effective, TTA necessitates access to training samples and modifies the training process.
-To surmount this challenge, TENT (Wang et al.
-2021) introduces the concept of fully test-time adaptation, utilizing only target data.
-TENT adapts the batch normalization layer by minimizing the entropy of model predictions.
-Distinct from TENT, recent works (Schneider et al.
-2020; Nado et al.
-2020) estimate the mean and variance of the activation in each batch normalization on the forthcoming test mini-batch.
-Consequently, LAME (Boudiaf et al.
-2022) and EATA (Niu et al.
-2022) concentrate on tackling the issue of catastrophic forgetting as the source model undergoes continuous updates.
-TSD (Wang, Zhang et al.
-2023) integrates self-training into TTA, selecting high-quality test samples to fit the model.
+  • Emerging research direction: We highlight the setting of the Few-Shot Test Time Adaptation (FS-TTA), where a limited few-shot support set is available prior to test-time adaptation.
+Leveraging these few-shot samples enables more effective mitigation of domain shift.
+  • Innovative framework: We propose a meticulously designed framework to address FS-TTA, incorporating fine-tuning of the pre-trained source model with a Feature Diversity Augmentation (FDA) module and performing test-time adaptation through high-quality pseudo-labeled samples in a self-training manner.
+  • State-of-the-Art Performance: Extensive empirical evaluations on multiple cross-domain classification benchmarks validate the effectiveness of our framework.
+Compared to the current state-of-the-art TTA methods, our approach achieves performance gains of 2.0% on PACS, 7.8% on OfficeHome, and 3.9% on DomainNet.
 Table 1: Comparison with various adaptation settings, where $s$ and $t$ denote source domain and target domain, respectively.
@@
 Source domain(s) Target domain Size of available target data
+Domain Generalization $L^{s_{1}},\dots,L^{s_{N}}$ 0
+Source-Free Domain Adaptation Pre-trained model on $L^{s_{1}},\dots,L^{s_{N}}$ Entire $U^{t}$ $\|U^{t}\|$
+Few-Shot Transfer Learning Pre-trained model on $L^{s_{1}},\dots,L^{s_{N}}$ Few-shot support set $L^{spt}\subset L^{t}$ $k\times\text{C}$
 […]
 […]
-[…]
-[…]
-[…]
-Comparisons with Other Settings
-We compare Few-Shot Test Time Adaptation (FS-TTA) with similar problem settings (details are in the appendix), as illustrated in Tab. 1.
-  • Compared with Domain Generalization, FS-TTA eliminates the need for access to source data, thereby safeguarding the privacy of the source data.
-Moreover, FS-TTA facilitates adaptation to the downstream target domain via model parameter updates.
-  • Compared with Source-Free Domain Adaptation, FS-TTA obviates the need to acquire all target domain data simultaneously and can dynamically update the model online, contingent upon incoming mini-batches.
-  • Compared with Few-Shot Transfer Learning, FS-TTA can not only utilize a small number of target samples, but can also further update the model using online mini-batch target data during the test time.
-  • Compared with Test Time Adaptation, FS-TTA capitalizes on an auxiliary small set of samples from the target, allowing the pre-trained source model to adapt more swiftly and proficiently to the target domain.
-Furthermore, FS-TTA demonstrates exceptional performance in managing circumstances involving substantial domain shifts.
-Method
-Problem Setting
-Considering a typical scenario where a source model $f_{\theta_{s}}$, is equipped with parameters $\theta_{s}$ and trained on source datasets $\mathcal{D}_{s_{1}},\mathcal{D}_{s_{2}},...,\mathcal{D}_{s_{n}}$, our goal is to adapt this pre-trained source model to a target domain $D_{t}$ without accessing source data.
-We postulate the existence of a small, labeled support set $S=\{(s_{i},y_{i})\}$ derived from $D_{t}$, where $s_{i}$ represents the image, and $y_{i}$ is its corresponding label.
-During the test time, mini-batches unlabeled target samples online arrive.
-Few-Shot Test Time Adaptation (FS-TTA) strives to effectively adapt the pre-trained source model $f_{\theta_{s}}$, utilizing the labeled support set $S$ and unlabeled online mini-batch to tackle the domain shift challenge.
-Notably, the support set $S$ could be offline obtained before the test time.
-Stage I: Model Adaptation via Fine-Tuning
-To significantly and swiftly enhance the initialization performance of the pre-trained source model in the target domain and minimize domain shifts, we design to fine-tune the pre-trained source model using the few-shot support set.
-Given the limited number of samples per class, there is a potential risk of overfitting during the fine-tuning process.
-To mitigate this, we introduce the Feature Diversity Augmentation (FDA) module, which generates new features by mixing statistics.
-Ultimately, we use a supervised classification loss to fine-tune the pre-trained source model.
-This entire procedure is illustrated in Stage I of Fig.
-3.
+2 Related Work
+2.1 Domain Generalization
+Domain Generalization (DG) aims to train models on multiple related but distinct source domains to ensure effective performance on unseen target domains.
+To enhance robustness, DG techniques often employ strategies such as data augmentation (Huang et al, 2021; Volpi et al, 2018) and data generation (Zhou et al, 2021; Robey et al, 2021) to introduce greater diversity during training.
+Other prevalent approaches leverage representation learning to extract domain-invariant features.
+This includes kernel-based methods (Li et al, 2018b) that project data into a shared feature space, domain adversarial learning (Sicilia et al, 2023) that aligns distributions via adversarial objectives, and invariant risk minimization (Krueger et al, 2021) which encourages models to perform consistently across domains.
+In addition, self-supervised (Kim et al, 2021) and meta-learning-based techniques (Chen et al, 2022) have been explored to further improve generalization.
+However, without exposure to the target domain, generalization remains inherently limited.
+2.2 Source-Free Domain Adaptation
+Source-Free Domain Adaptation (SFDA) aims to adapt a pre-trained source model to an unlabeled target domain while ensuring that no source data is accessed during the adaptation process.
+By eliminating the dependence on source data, SFDA effectively safeguards source data privacy, making it particularly suitable for scenarios where data sharing is restricted.
+SFDA techniques can be broadly categorized into two main approaches: pseudo-label strategies and generative methods.
+The former leverages target pseudo-labels to facilitate self-training, thereby enabling implicit adaptation without requiring explicit supervision (Tanwisuth et al, 2021; Ahmed et al, 2021; Liang et al, 2021; Xin et al, 2023).
+The latter employs generative models to synthesize target-style training data, allowing the model to bridge the domain gap through data augmentation and distribution alignment (Qiu et al, 2021; Liu et al, 2021).
+Similar to SFDA, our proposed Few-Shot Test Time Adaptation (FS-TTA) also maintains the source-free property, ensuring that adaptation is performed without relying on source data while leveraging a small support set to enhance adaptation efficiency.
+2.3 Few-Shot Transfer Learning
+Test Time Adaptation (TTA) aims to adapt a pre-trained source model on-the-fly during inference to mitigate distribution shifts.
+Early TTA methods apply self-supervised learning objectives (Sun et al, 2020), but typically require access to training data or modification of the training process.
+TENT (Wang et al, 2021) addresses this by proposing fully test-time adaptation, relying solely on target samples and adapting batch normalization parameters via entropy minimization.
+Subsequent approaches such as (Schneider et al, 2020; Nado et al, 2020) update statistics on each incoming mini-batch, while methods like LAME (Boudiaf et al, 2022) and EATA (Niu et al, 2022) tackle catastrophic forgetting during continual adaptation.
+TSD (Wang et al, 2023) further integrates self-training to selectively update using confident predictions.
+As a result, adaptation often relies heavily on the quality of incoming test samples.
+2.4 Test Time Adaptation
+Test Time Adaptation (TTA) aims to adapt a pre-trained source model during inference to mitigate distribution shifts between training and test domains.
+Early TTA methods address this challenge through self-supervised auxiliary tasks (Sun et al, 2020), which, while effective, often require access to training data or modifications to the training procedure.
+To overcome this limitation, TENT (Wang et al, 2021) proposes fully test-time adaptation by leveraging only target data, updating batch normalization parameters via entropy minimization.
+Building on this, subsequent works (Schneider et al, 2020; Nado et al, 2020) estimate batch normalization statistics dynamically from incoming test batches.
+Other approaches, such as LAME (Boudiaf et al, 2022) and EATA (Niu et al, 2022), focus on preventing catastrophic forgetting during continuous model updates.
+More recently, TSD (Wang et al, 2023) incorporates self-training by selecting confident test samples to guide adaptation.
+Despite their progress, these methods heavily rely on the quality and stability of online target data.
+2.5 Comparisons with Other Settings
+We compare Few-Shot Test Time Adaptation (FS-TTA) with similar problem settings (details are in the appendix), as illustrated in Table 1.
+  • Compared with Domain Generalization, FS-TTA eliminates the necessity of accessing source data, thereby ensuring the preservation of source data privacy.
+Moreover, it allows for adaptation to the downstream target domain by updating model parameters, making it more flexible and applicable in real-world settings.
+  • Compared with Source-Free Domain Adaptation, FS-TTA removes the constraint of requiring all target domain data to be available at once.
+Instead, it facilitates dynamic and continuous online model updates, enabling adaptation based on incoming mini-batches of target data, which is particularly beneficial in streaming or real-time applications.
+  • Compared with Few-Shot Transfer Learning, FS-TTA not only makes use of a limited number of target domain samples for adaptation but also continuously refines the model during test time by incorporating online mini-batch target data.
+This ensures more efficient and progressive adaptation to changing target distributions.
+  • Compared with Test Time Adaptation, FS-TTA leverages a small auxiliary set of target samples, allowing the pre-trained source model to adapt more quickly and effectively to the target domain.
+Additionally, FS-TTA demonstrates superior performance in handling challenging scenarios where there are substantial domain shifts, making it a more robust and reliable solution.
+3 Preliminary
+3.1 Instance Normalization
+Instance Normalization (Ulyanov et al, 2016) is a normalization technique widely used in deep neural network architectures, especially in the context of style transfer and generative models.
+Let us consider a batch images with size $N\times C\times H\times W$, where $N$ is the batch size, $C$ is the number of channels, and $H$ and $W$ are the height and width of the images.
+For each sample $i$ and channel $c$, we compute the mean and standard deviation as follows:
+$\mu_{i,c}=\frac{1}{H\times W}\sum_{h=1}^{H}\sum_{w=1}^{W}x_{i,c,h,w},$ (1)
+$\sigma_{i,c}=\sqrt{\frac{1}{H\times W}\sum_{h=1}^{H}\sum_{w=1}^{W}(x_{i,c,h,w}-\mu_{i,c})^{2}},$ (2)
+where $x_{i,c,h,w}$ denotes the input feature of samples $i$, channel $c$, height $h$, and width $w$. After computing the mean and standard deviation, we can normalize the input features:
+${\rm IN}(x_{i,c,h,w})=\gamma\frac{x_{i,c,h,w}-\mu_{i,c}}{\sigma_{i,c}}+\beta,$ (3)
+where $\gamma,\beta\in R^{C}$ are learnable transformation parameters.
+3.2 Class Prototype
+The class prototype is a representative point in the feature space that summarizes the key characteristics of a class.
+For each class, it serves as a centroid or an anchor point around which samples of the class cluster.
+Let us denote $F=\{f_{1},f_{2},...,f_{n}\}$ as a set of $n$ sample features in class $c$, where each $f_{i}\in R^{d}$ represents a $d$-dimensional feature vector of a sample.
+The prototype $P$ of the class $c$ is calculated as the mean of all feature vectors, namely that:
+$P_{c}=\frac{1}{n}\sum_{i=1}^{n}f_{i},$ (4)
+where $P_{c}\in R^{d}$. The class prototype plays an important role in few-shot scenarios.
 Figure 3: Illustration of our two-stage framework.
@@
 In order to update the prototype memory bank and model with effective samples, we propose the entropy filter and consistency selection modules.
-Feature Diversity Augmentation.
-Prior research (Zhou et al.
-2021) has demonstrated a significant association between feature statistics and image style, which is intricately linked to data distribution within the field of computer vision.
-Inspired by their discoveries, we propose Feature Diversity Augmentation (FDA), which enhances feature diversity and mitigates the risk of overfitting during fine-tuning.
-FDA is incorporated between layers (blocks) in the pre-trained source backbone, as depicted in Fig.
-3.
+4 Method
+4.1 Problem Setting
+Considering a typical scenario where a source model $f_{\theta_{s}}$ is equipped with parameters $\theta_{s}$ and trained on source datasets $\mathcal{D}{s_{1}},\mathcal{D}{s_{2}},…,\mathcal{D}{s_{n}}$, our objective is to adapt this pre-trained model to a target domain $D_{t}$ without accessing source data.
+A small, labeled support set $S={(s_{i},y_{i})}$ is provided from $D_{t}$, where $s_{i}$ denotes an image and $y_{i}$ its corresponding label.
+During test time, unlabeled target samples arrive sequentially in mini-batches.
+Few-Shot Test Time Adaptation (FS-TTA) aims to effectively adapt the source model $f_{\theta_{s}}$, by leveraging the support set $S$ in conjunction with the streaming unlabeled data to mitigate domain shift.
+Notably, the support set $S$ can be acquired offline prior to deployment, and in many real-world applications, collecting such limited supervision is both feasible and cost-efficient.
+4.2 Stage I: Model Adaptation via Fine-Tuning
+To significantly and swiftly enhance the initialization performance of the pre-trained source model in the target domain and minimize domain shifts, we design to fine-tune the pre-trained source model using the few-shot support set.
+Given the limited number of samples per class, there is a potential risk of overfitting during the fine-tuning process.
+To mitigate this, we introduce the Feature Diversity Augmentation (FDA) module, which generates new features by mixing statistics.
+Ultimately, we use a supervised classification loss to fine-tune the pre-trained source model.
+This entire procedure is illustrated in Stage I of Figure 3.
+Feature Diversity Augmentation (FDA).
+Prior research (Zhou et al, 2021) has demonstrated a significant association between feature statistics and image style, which is intricately linked to data distribution within the field of computer vision.
+To increase style diversity while preserving semantic consistency, we introduce Feature Diversity Augmentation (FDA), a feature-level data augmentation technique that simulates various image styles without altering the original class labels.
+This approach effectively enriches the support set and helps reduce the risk of overfitting during fine-tuning.
+FDA is incorporated between layers (blocks) in the pre-trained source backbone, as depicted in Figure 3.
 More specifically, FDA mixes the feature statistics of two random samples to generate new features.
-The computations within FDA module can be summarized in three steps.
+The computations within the FDA module can be summarized in three steps.
 Firstly, given two feature maps $f_{i}$ and $f_{j}$ from the support set, we compute their feature statistics ($\mu_{i},\sigma_{i}$) and ($\mu_{j},\sigma_{j}$).
 Secondly, FDA generates the mixtures of feature statistics:
-$\boldsymbol{\gamma}_{\text{mix }}=\lambda\sigma_{i}+(1-\lambda)\sigma_{j},$ (1)
-$\boldsymbol{\beta}_{\text{mix }}=\lambda\mu_{i}+(1-\lambda)\mu_{j}.$ (2)
+$\bm{\gamma}_{\text{mix }}=\lambda\sigma_{i}+(1-\lambda)\sigma_{j},$ (5)
+$\bm{\beta}_{\text{mix }}=\lambda\mu_{i}+(1-\lambda)\mu_{j}.$ (6)
 In this case, $\lambda$ denotes the mixing ratio coefficient. Ultimately, the mixtures of feature statistics are applied to the feature map $f_{i}$ via instance normalization:
-$f_{i}^{\prime}=\boldsymbol{\gamma}_{\operatorname{mix}}\odot\frac{f_{i}-\mu_{i}}{\sigma_{i}}+\boldsymbol{\beta}_{\operatorname{mix}},$ (3)
+$f_{i}^{\prime}=\bm{\gamma}_{\operatorname{mix}}\odot\frac{f_{i}-\mu_{i}}{\sigma_{i}}+\bm{\beta}_{\operatorname{mix}},$ (7)
 where $f_{i}^{\prime}$ represents the newly generated feature map.
@@
 Specifically, the few-shot support set is processed through $f_{\theta_{s}}$ to minimize a supervised loss, defined as:
-$\mathcal{L}_{\text{cls}}=-\sum_{i=1}^{k*C}\mathcal{H}\left(y_{i},p\left(\hat{y_{i}}\mid\boldsymbol{s}_{i}\right)\right),$ (4)
+$\mathcal{L}_{\text{cls}}=-\sum_{i=1}^{k*C}\mathcal{H}\left(y_{i},p\left(\hat{y_{i}}\mid\bm{s}_{i}\right)\right),$ (8)
 where $\mathcal{H}(\cdot)$ is the cross-entropy loss.
 The term $y_{i}$ is the ground-truth label of $s_{i}$, indicating one of sample from few-shot support set, and $C$ represents categories of the target.
-Table 2: Compared with test time adaptation methods on three datasets with ResNet50 backbone.
-OfficeHome PACS DomainNet
-Method Art Clip Prod Real Avg. Art Cart Phot Sket Avg. Avg.
-Test time adaptation methods
-ERM (Vapnik 1999) 60.7 55.7 76.2 76.8 67.4 82.5 80.8 94.0 80.9 84.5 45.2
-BN (Nado et al. 2020) 58.2 55.6 75.1 75.5 66.1 83.2 84.9 94.0 77.9 85.0 43.3
-TENT (Wang et al. 2021) 60.6 58.7 76.5 76.1 68.0 85.2 86.7 94.9 82.9 87.4 44.7
-T3A (Iwasawa and Matsuo 2021) 61.2 56.7 78.0 77.3 68.3 84.0 82.3 95.0 82.7 86.0 46.1
-ETA (Niu et al. 2022) 58.4 55.8 75.2 75.5 66.2 83.2 84.9 94.0 77.9 85.0 46.1
-LAME (Boudiaf et al. 2022) 58.7 55.6 75.1 75.4 66.2 84.9 85.5 95.0 80.9 86.6 43.2
-TSD (Wang, Zhang et al. 2023) 62.3 57.5 77.5 77.5 68.7 87.6 88.7 96.1 85.0 89.4 47.7
-Fine-tuning + Test time adaptation methods
-FT+TENT (Wang et al. 2021) 68.8 65.5 79.8 78.5 73.2 87.0 86.9 95.2 83.6 88.2 45.4
-FT+TSD (Wang, Zhang et al. 2023) 70.5 65.1 80.3 79.2 73.8 88.3 88.6 96.5 85.9 89.8 48.5
-FS-TTA 73.2 68.3 83.0 81.6 76.5 90.4 89.7 97.6 87.8 91.4 51.6
-$\Delta_{up}$ over TSD (+10.9$\uparrow$ (+10.8$\uparrow$ (+5.5$\uparrow$ (+4.1$\uparrow$ (+7.8$\uparrow$ (+2.8$\uparrow$ (+1.0$\uparrow$ (+1.5$\uparrow$ (+2.8$\uparrow$ (+2.0$\uparrow$ (+3.9$\uparrow$
-Stage II: Test Time Adaptation
+4.3 Stage II: Test Time Adaptation
 During this stage, a mini-batch of unlabeled samples, denoted as $x=\{x_{1},x_{2},..,x_{B}\}$, online arrives.
@@
 To address this issue, we propose two modules to produce high quality pseudo-labels.
-The first is entropy filter, which screens out unreliable samples using Shannon entropy (Shannon 2001).
+The first is entropy filter, which screens out unreliable samples using Shannon entropy (Shannon, 2001).
 Typically, samples with higher entropy are considered to have lower prediction confidence.
@@
 After that, pseudo-labels with consistency prediction is preserved for model adaptation.
-The entire process is outlined in Stage II of Fig.
-3.
+The entire process is outlined in Stage II of Figure 3.
 Entropy Filter.
 To dynamically update the model using online mini-batch target, it is crucial to filter out noisy samples, as they may be assigned to incorrect classes, resulting in inaccurate prototype computation.
-In this regard, we propose the Entropy Filter, which employs Shannon entropy (Shannon 2001) to select confident samples in the mini-batch.
+In this regard, we propose the Entropy Filter, which employs Shannon entropy (Shannon, 2001) to select confident samples in the mini-batch.
 For an sample $x_{i}$, its entropy can be computed as:
 $H(p_{i})=-\sum(p_{i})\cdot log(p_{i}).$ (9)
-Based on the insights from previous work (Wang et al.
-2021), high entropy samples should be filtered out, as lower entropy typically indicates higher accuracy.
-Consequently, we sort the entropy of all samples in the mini-batch and select the top $\alpha\%$ samples with lower entropy, donated as $\hat{x}=\{\hat{x}_{1},\hat{x}_{2},...,\hat{x}_{|\alpha\%*B|}\}$.
+Based on the insights from previous work (Wang et al, 2021), high entropy samples should be filtered out, as lower entropy typically indicates higher accuracy.
+Consequently, we sort the entropy of all samples in the mini-batch and select the top $\alpha\%$ samples with lower entropy, donated as $\hat{x}=\{\hat{x}_{1},\hat{x}_{2},\ldots,\hat{x}_{\lfloor\alpha\cdot B\rfloor}\}$.
 Prototype Memory Bank.
@@
 The prototype memory bank is initialized with the few-shot support set $S$, defined as:
-$m_{c_{0}}=\frac{\sum_{i=1}^{|S|}f_{i}\cdot\mathbbm{1}[y_{i}=c]}{\sum_{i=1}^{|S|}\mathbbm{1}[y_{i}=c]},$ (6)
-where $\mathbbm{1}[\cdot]$represents an indicator function, yielding a value of 1 if the argument is true or 0 otherwise, and $m_{c_{0}}$ denotes the initial moment of the $c$-th class prototype.
+$m_{c_{0}}=\frac{\sum_{i=1}^{\lvert S\rvert}f_{i}\cdot\mathds{1}[y_{i}=c]}{\sum_{i=1}^{\lvert S\rvert}\mathds{1}[y_{i}=c]},$ (10)
+where $\mathds{1}[\cdot]$ represents an indicator function, yielding a value of 1 if the argument is true or 0 otherwise, and $m_{c_{0}}$ denotes the initial moment of the $c$-th class prototype.
 Thanks to the few-shot support, precise guidance can be provided during the initial phase, thereby reducing reliance on the quality of online mini-batch data.
 Throughout the test time adaptation process, we persistently update the prototype memory bank by incorporating selected reliable samples with pseudo labels:
-$m_{c_{t}}=\beta*m_{c_{t-1}}+(1-\beta)*\frac{\sum_{j=1}^{|\hat{x}|}f_{j}\cdot\mathbbm{1}[\hat{y}_{j}=c]}{\sum_{j=1}^{|\hat{x}|}\mathbbm{1}[\hat{y}_{j}=c]},$ (7)
+$m_{c_{t}}=\beta\cdot m_{c_{t-1}}+(1-\beta)\cdot\frac{\sum_{j=1}^{\lvert\hat{x}\rvert}f_{j}\cdot\mathds{1}[\hat{y}_{j}=c]}{\sum_{j=1}^{\lvert\hat{x}\rvert}\mathds{1}[\hat{y}_{j}=c]}.$ (11)
 where $m_{c_{t}}$ represents the $c$-th class prototype at time $t$, and $\beta$ represents the sliding update coefficient.
+Table 2: Table 2: Comparison with test-time adaptation methods on three datasets with ResNet-50 backbone.
+FS-TTA achieves consistent improvements over TSD (Wang et al, 2021), the strongest baseline method.
+OfficeHome PACS DomainNet
+Method Art Clip Prod Real Avg. Art Cart Phot Sket Avg. Avg.
+Test time adaptation methods
+ERM (Vapnik, 1999) 60.7 55.7 76.2 76.8 67.4 82.5 80.8 94.0 80.9 84.5 45.2
+BN (Nado et al, 2020) 58.2 55.6 75.1 75.5 66.1 83.2 84.9 94.0 77.9 85.0 43.3
+TENT (Wang et al, 2021) 60.6 58.7 76.5 76.1 68.0 85.2 86.7 94.9 82.9 87.4 44.7
+T3A (Iwasawa and Matsuo, 2021) 61.2 56.7 78.0 77.3 68.3 84.0 82.3 95.0 82.7 86.0 46.1
+ETA (Niu et al, 2022) 58.4 55.8 75.2 75.5 66.2 83.2 84.9 94.0 77.9 85.0 46.1
+LAME (Boudiaf et al, 2022) 58.7 55.6 75.1 75.4 66.2 84.9 85.5 95.0 80.9 86.6 43.2
+TSD (Wang et al, 2023) 62.3 57.5 77.5 77.5 68.7 87.6 88.7 96.1 85.0 89.4 47.7
+PROGRAM (Sun et al, 2024) 63.4 54.3 77.2 77.2 68.0 87.2 84.1 96.9 76.4 86.2 43.3
+DEYO (Lee et al, 2024) 63.8 54.9 76.4 77.3 68.1 88.4 85.2 97.1 82.3 88.2 42.5
+Fine-tuning + Test time adaptation methods
+FT+TENT (Wang et al, 2021) 68.8 65.5 79.8 78.5 73.2 87.0 86.9 95.2 83.6 88.2 45.4
+FT+TSD (Wang et al, 2023) 70.5 65.1 80.3 79.2 73.8 88.3 88.6 96.5 85.9 89.8 48.5
+FS-TTA 73.2 68.3 83.0 81.6 76.5 90.4 89.7 97.6 87.8 91.4 51.6
+$\Delta_{up}$ over TSD (+10.9$\uparrow$ (+10.8$\uparrow$ (+5.5$\uparrow$ (+4.1$\uparrow$ (+7.8$\uparrow$ (+2.8$\uparrow$ (+1.0$\uparrow$ (+1.5$\uparrow$ (+2.8$\uparrow$ (+2.0$\uparrow$ (+3.9$\uparrow$
 Test Time Adaptation.
@@
 This strategy can be implemented through a filter mask for samples $x_{j}$ as follows:
-$\mathcal{M}_{j}=\mathbbm{1}\left[\arg\max p_{j}=\arg\max\hat{p}_{j}\right].$ (9)
+$\mathcal{M}_{j}=\mathbbm{1}[\arg\max p_{j}=\arg\max\hat{p}_{j}].$ (13)
 Ultimately, we can update the model using reliable samples, and the loss can be formulated as follows:
-$\mathcal{L}_{online}=\frac{{\textstyle\sum_{j=1}^{|\hat{x}|}}\mathcal{H}_{j}*\mathcal{M}_{j}}{{\textstyle\sum_{j=1}^{|\hat{x}|}}\mathcal{M}_{j}}.$ (10)
+$\mathcal{L}_{online}=\frac{{\textstyle\sum_{j=1}^{\|\hat{x}\|}}\mathcal{H}_{j}*\mathcal{M}_{j}}{{\textstyle\sum_{j=1}^{\|\hat{x}\|}}\mathcal{M}_{j}}.$ (14)
 It’s noteworthy that our self-training process does not involve specifying any threshold, which enhances the model’s generalizability.
-Experiment
-Experimental Settings
-Dataset.
-To test the effectiveness of our setting and method, we experiment on three cross-domain benchmarks.
-PACS (Li et al.
-2017) consists of four distinct domains: Art, Cartoon, Photo, and Sketch, and comprises a total of 9,991 images and 7 classes.
-Office-Home (Venkateswara et al.
-2017) consists of 15,588 images from four distinct image domains, namely Real World, Clipart, Art, and Product, with 65 classes.
-DomainNet (Peng et al.
-2019) is a large-scale dataset that comprises six domains, denoted as $d\in${Clipart, Infograph, Painting, Quickdraw, Real, Sketch}, with a total of 586,575 images and 345 classes.
+5 Experiment
+5.1 Experimental Settings
+Dataset. To evaluate the effectiveness of our proposed setting and method, we conduct experiments on three cross-domain benchmarks.
+  • PACS (Li et al, 2017) consists of 9,991 images spanning four distinct domains: Art, Cartoon (Cart), Photo (Phot), and Sketch (Sket).
+Each domain contains seven object categories: dog, elephant, giraffe, guitar, horse, house, and person.
+  • Office-Home (Venkateswara et al, 2017) comprises 15,588 images distributed across four domains: Art, Clipart (Clip), Product (Prod), and Real-World (Real), with each domain encompassing 65 image categories.
+  • DomainNet (Peng et al, 2019) is a large-scale dataset containing six domains: Clipart (Clip), Infograph (Info), Painting (Pain), Quickdraw (Quic), Real, Sketch (Sket), comprising a total of 586,575 images across 345 classes.
 Implementation Details.
-In the main experiments, we utilize ResNet-50 (He et al.
-2016) pre-trained on ImageNet-1k (Russakovsky et al.
-2015) as our backbone, which is widely used in test time adaptation literature.
-For source model training, we adhere to the leave-one-domain-out protocol, as recommended by previous work  (Wang, Zhang et al.
-2023; Zhou et al.
-2021).
-We employ the Adam optimizer with a learning rate of $5e^{-5}$.
+In our main experiments, we employ ResNet-50(He et al, 2016), pre-trained on ImageNet-1k(Russakovsky et al, 2015), as the backbone model, as it is widely adopted in the test-time adaptation literature.
+For source model training, we follow the leave-one-domain-out protocol, as recommended by prior studies (Wang et al, 2023; Zhou et al, 2021), treating one domain as the unlabeled target and the rest as source domains.We set the batch size to 32 for each source domain and use a learning rate of 5e-5.
+Both the dropout probability and weight decay are set to zero.
+The source model is trained for 5,000 iterations, except for DomainNet, where we extend training to 15,000 iterations, following the methodology in (Cha et al, 2021).
+All images are resized to 224 × 224, and data augmentation is applied during source domain training, including random cropping, horizontal flipping, color jittering, and intensity adjustments.
 For few-shot test time adaptation, we also employ the Adam optimizer and set the batch size to The few-shot support set typically selects 5 to 16 samples per class, depending on the difficulty of the target.
 We carry out all experiments on NVIDIA V100 GPUs.
-Baselines.
-We compare our method with test time adaptation methods, BN (Nado et al.
-2020), TENT (Wang et al.
-2021), ETA (Niu et al.
-2022), T3A (Iwasawa and Matsuo 2021), LAME (Boudiaf et al.
-2022), and TSD (Wang, Zhang et al.
-2023).
-Additionally, we create new baselines by combining fine-tuning with TTA methods for a more comprehensive comparison.
-We also compare with some methods in domain generalization and source-free domain adaptation, including DNA (Chu et al.
-2022), PCL (Yao et al.
-2022), SWAD (Cha et al.
-2021), and F-mix (Kundu et al.
-2022).
+Figure 4: Comprehensive comparison between our method and the state-of-the-art method in DG/TTA settings on DomainNet.
 Table 3: Compared with existing DG and SFDA methods on OfficeHome and DomainNet.
@@
 Domain generalization methods
-ERM (Vapnik 1999) 60.7 55.7 76.2 76.8 67.4 64.8 22.1 51.8 13.8 64.7 54.0 45.2
-DNA (Chu et al. 2022) 67.7 57.7 78.9 80.5 71.2 66.1 23.0 54.6 16.7 65.8 56.8 47.2
-PCL (Yao et al. 2022) 67.3 59.9 78.7 80.7 71.6 67.9 24.3 55.3 15.7 66.6 56.4 47.7
-SWAD (Cha et al. 2021) 66.1 57.7 78.4 80.2 70.6 66.1 22.4 53.6 16.3 65.5 56.2 46.7
+ERM (Vapnik, 1999) 60.7 55.7 76.2 76.8 67.4 64.8 22.1 51.8 13.8 64.7 54.0 45.2
+DNA (Chu et al, 2022) 67.7 57.7 78.9 80.5 71.2 66.1 23.0 54.6 16.7 65.8 56.8 47.2
+PCL (Yao et al, 2022) 67.3 59.9 78.7 80.7 71.6 67.9 24.3 55.3 15.7 66.6 56.4 47.7
+SWAD (Cha et al, 2021) 66.1 57.7 78.4 80.2 70.6 66.1 22.4 53.6 16.3 65.5 56.2 46.7
 Source-free domain adaptation methods
-F-mix (Kundu et al. 2022) 72.6 67.4 85.9 83.6 77.4 75.4 24.6 57.8 23.6 65.8 58.5 51.0
+F-mix (Kundu et al, 2022) 72.6 67.4 85.9 83.6 77.4 75.4 24.6 57.8 23.6 65.8 58.5 51.0
 FS-TTA 73.2 68.3 83.0 81.6 76.5 68.6 30.8 56.4 24.2 69.1 60.2 51.6
 SWAD + FS-TTA 77.4 71.1 86.4 84.2 79.8 - - - - - - -
-Performance Comparisons
+Baselines.
+We compare our method with various test-time adaptation (TTA) approaches, including BN(Nado et al, 2020), TENT(Wang et al, 2021), ETA(Niu et al, 2022), T3A(Iwasawa and Matsuo, 2021), LAME(Boudiaf et al, 2022), TSD(Wang et al, 2023),PROGRAM (Sun et al, 2024) and DEYO (Lee et al, 2024).
+Additionally, we establish new baselines by integrating fine-tuning with existing TTA methods to ensure a more comprehensive comparison.
+Furthermore, we compare our approach with selected methods from domain generalization, source-free domain adaptation, including DNA(Chu et al, 2022), PCL(Yao et al, 2022), SWAD(Cha et al, 2021), and F-mix(Kundu et al, 2022).
+Finally, we set up a comparison with the few-shot transfer learning methods, including AdaBN (Li et al, 2016), $L^{2}$ (Li et al, 2018a), DELTA (Li et al, 2019), FLUTE (Triantafillou et al, 2021), LCCS (Zhang et al, 2022).
+For a global overview, we compare our method with state-of-the-art method in various settings, as shown in Figure 4.
+5.2 Performance Comparisons
 Comparison with TTA methods.
-Tab.
-2 details the comparison results between our method and various TTA methods on the Office-Home and PACS datasets, as well as the final results of DomainNet (detailed in Tab.
-3).
+Table 2 details the comparison results between our method and various TTA methods on the Office-Home and PACS datasets, as well as the final results of DomainNet (detailed in Table 3).
 We observe that our method achieves state-of-the-art performance.
@@
 On the other two datasets, we observe average performance increments of 6.9% (PACS) and 6.4% (DomainNet).
+Table 4: Compared with few-shot transfer learning methods on PACS dataset.
+PACS
+Method Art Cart Phot Sket Avg.
+Few-shot transfer learning methods
+AdaBN (Li et al, 2016) 85.0 83.5 96.0 78.7 85.8
+$L^{2}$ (Li et al, 2018a) 85.6 84.1 96.4 76.3 85.6
+DELTA (Li et al, 2019) 85.6 83.8 96.5 76.3 85.6
+FLUTE (Triantafillou et al, 2021) 87.2 86.1 97.2 81.7 88.1
+LCCS (Zhang et al, 2022) 87.7 86.9 97.5 83.0 88.8
+FS-TTA 90.4 89.8 97.6 87.9 91.4
 Moreover, our method outperforms the state-of-the-art TTA method, TSD, with average performance increments of 2.0% (PACS), 7.8% (Office-Home), and 3.9% (DomainNet).
@@
 In conclusion, our FS-TTA demonstrates a notable advantage in tasks that closely resemble real-world scenarios and provides a significant boost in performance with minimal additional computational overhead.
-Finally, for a more comprehensive comparison with TTA methods, we construct new baselines, which combine fine-tuning with TTA methods (e.g., TENT and TSD).
-According to the results in Tab.
-2, our method shows an average (three datasets) improvement of 4.2% over Fine-Tuning+TENT, and 2.4% over Fine-Tuning+TSD, which demonstrates the superiority of our method over existing TTA techniques in migrating to few-shot TTA scenarios.
-This result benefits from our unique FDA module design in Fine-Tuning and the use of few-shot support set to initialize prototype memory bank.
-Refer to captionRefer to caption(a) Different stage(b) Different $\alpha$
-Figure 4: (a) effectiveness analysis about two-stage framework and (b) sensitivity analysis about parameter $\alpha$.
+Finally, for a more comprehensive comparison with TTA methods, we construct new baselines by combining fine-tuning with representative TTA approaches.
+Specifically, we select TENT (Wang et al, 2021), as a widely adopted and foundational method in test-time adaptation, and TSD (Wang et al, 2023), which demonstrates state-of-the-art performance across benchmarks.
+According to the results in Table 2, our method achieves an average improvement of 4.2% over Fine-Tuning+TENT and 2.4% over Fine-Tuning+TSD across the three datasets.
+These results highlight the superiority of our framework in migrating to the few-shot TTA setting, benefiting from the proposed FDA module and the support-set-based prototype initialization.
+Figure 5: Ablation Study on (a) effectiveness analysis about two-stage framework and (b) effectiveness analysis about FDA module.
+Figure 6: Comprehensive Sensitivity Analysis of (1) Different Entropy Filter Proportion $\alpha$ and K-shot Number on Model Performance.
 Comparison with DG/SFDA methods.
@@
 To answer this question, we compare our method with several methods in DG and SFDA.
-The results of Office-Home dataset are shown in Tab.
-3.
+The results of Office-Home dataset are shown in Table 3.
 It can be seen that our method outperforms the state-of-the-art methods in DG, such as SWAD and PCL.
@@
 It is worth noting that FS-TTA is more flexible in real-world scenarios than SFDA since it adapts the target data in an offline manner, requiring more training loops and resources.
-The results of DomainNet are shown in Tab.
-3.
+The results of DomainNet are shown in Table 3.
 The overall performance of FS-TTA outperforms the SFDA methods, suggesting that FS-TTA is more adept at handling challenging tasks.
-Ablation Study
-Figure 5: The t-SNE feature visualization of (a) ERM, (b) TENT, (c) LAME, and (d) our method.
+Comparison with few-shot transfer learning methods.
+In our research, we focus on Few-Shot Test Time Adaptation (FS-TTA), which utilizes a small number of target domain samples to enhance adaptation.
+To ensure a comprehensive evaluation, we compare our approach with existing few-shot transfer learning methods.
+The results on the PACS dataset are presented in Table 4.
+According to the results, FS-TTA consistently outperforms all baseline methods across different domains, achieving the highest average accuracy of 91.4%, which surpasses the best-performing baseline LCCS (88.8%).
+This result highlights the effectiveness of our approach in adapting to domain shifts and improving classification performance in the few-shot setting.
+5.3 Ablation Study
 Effectiveness of two-stage framework.
-Our proposed method consists of two stages, with the individual contributions of each stage presented in Fig.
-4(a).
+Our proposed method consists of two stages, with the individual contributions of each stage presented in Figure 5(a).
 Compared to the baseline source model, Stage I of our approach achieves an average improvement of 6.6% on the Office-Home.
@@
 As a result, our two-stage framework establishes itself as a robust foundation for the Few-Shot Test Time Adaptation setting, demonstrating its considerable potential in enabling online model adaptation in real-world situations where labeled data is scarce.
+Figure 7: The t-SNE feature visualization of (a) ERM, (b) TENT, (c) LAME, and (d) FS-TTA (Ours).
+Figure 8: Analyzing the Impact of Batch Size on Accuracy and Running Efficiency.
+Effectiveness of FDA module.
+In our first phase, we introduce the FDA module to tackle overfitting issues through feature augmentation.
+Here we conduct additional ablation experiments on the FDA module and compare it with Mix-up augmentation, as depicted in Figure 5(b).
+The results from the ablation experiments indicate that the FDA module is effective and outperforms mix-up augmentation.
+The baseline method (without any techniques in the fine-tuning phase) achieves an accuracy of 76.12%, while incorporating Mix-up leads to a slight improvement, reaching 76.2% (+0.08%).
+However, when our FDA module is introduced, the performance further increases to 76.49%, yielding a notable improvement of +0.37% over the baseline.
+These results highlight the advantage of FDA module in enhancing feature diversity and robustness, surpassing standard augmentation techniques like Mix-up.
 Sensitivity to $\alpha$.
@@
 To evaluate the impact of $\alpha$, we conduct an experimental analysis on the Office-Home dataset by assigning $\alpha$ to 0, 0.3, 0.6, and 1, respectively.
-The results, as shown in Fig.
-4(b), demonstrate that $\alpha>0$ yields performance improvements compared to $\alpha=0$ (the source model), highlighting the effectiveness of our proposed framework.
+The results, as shown in Figure 6(b), demonstrate that $\alpha>0$ yields performance improvements compared to $\alpha=0$ (the source model), highlighting the effectiveness of our proposed framework.
 Furthermore, $\alpha=0.3$ and $\alpha=0.6$ perform better than $\alpha=1$ (no filter), indicating the effectiveness of our entropy filter strategy.
-Qualitative analysis by t-SNE visualization.
-We present t-SNE visualizations to compare the feature representations of the pre-trained source model (ERM), test time adaptation methods (TENT and LAME), and our proposed method, as illustrated in Fig.
-5.
-The learned features of the pre-trained source model on the target domain are not well-separated due to the significant domain gap, as shown in Fig.
-5(a).
-Additionally, we can observe no considerable feature distribution changes on the target domain after adaptation with TENT and LAME methods, as shown in Fig.
-5(b) and Fig.
-5(c).
-In contrast, our method produces more uniform and aligned feature distribution after adapting to the target domain, as shown in Fig.
-5(d).
-Efficiency analysis.
-In our main experiments, we opt for a mini-batch size of 64.
-To examine the variations in performance and computational efficiency with different batch size during test-time adaptation, we conduct a series of analytical experiments.
-As shown in Fig.
-6(a), we observe that accuracy experiences a gradual increase as the batch size incrementally grows, reaching a plateau around a batch size of 64.
-In contrast, as shown in Fig.
-6(b), running time exhibits a decreasing trend as the batch size grows.
-However, beyond a batch size of 64, the running time appears to stabilize.
-Consequently, for real-world applications aiming to achieve a trade-off between accuracy and computational efficiency, we suggest a batch size in the vicinity of 64.
-Refer to captionRefer to caption(a) Accuracy (batch size)(b) Running Time
-Figure 6: During the test time adaptation, we analyze:(a) accuracy and (b) running time, across various batchsize.
-Figure 7: Ablation experiments on shot size.
 Ablation experiments on shot size.
@@
 For instance, the 3-shot configuration achieves a 3.8% performance improvement over the TSD.
-Conclusion
+Qualitative analysis by t-SNE visualization.
+We present t-SNE visualizations to compare the feature representations of the pre-trained source model (ERM), test time adaptation methods (TENT and LAME), and our proposed method, as illustrated in Figure 7.
+The learned features of the pre-trained source model on the target domain are not well-separated due to the significant domain gap, as shown in Figure 7(a).
+Additionally, we can observe no considerable feature distribution changes on the target domain after adaptation with TENT and LAME methods, as shown in Figure 7(b) and Figure 7(c).
+In contrast, our method produces more uniform and aligned feature distribution after adapting to the target domain, as shown in Figure 7(d).
+Efficiency analysis.
+In our main experiments, we opt for a mini-batch size of 64.
+To examine the variations in performance and computational efficiency with different batch size during test-time adaptation, we conduct a series of analytical experiments.
+As shown in Figure 8(a), we observe that accuracy experiences a gradual increase as the batch size incrementally grows, reaching a plateau around a batch size of 64.
+In contrast, as shown in Figure 8(b), running time exhibits a decreasing trend as the batch size grows.
+However, beyond a batch size of 64, the running time appears to stabilize.
+Consequently, for real-world applications aiming to achieve a trade-off between accuracy and computational efficiency, we suggest a batch size in the vicinity of 64.
+Table 5: Comparison of our method with the baseline TSD on both ResNet and ViT-B/16 backbones across the PACS and Office-Home datasets.
+Backbones PACS Office-Home
+ResNet 84.59 67.37
++ TSD (Wang et al, 2023) 89.41 68.67
++ Ours 91.42 76.49
+ViT-B/16 87.13 79.06
++ TSD (Wang et al, 2023) 90.20 81.80
++ Ours 91.89 87.32
+Scalability on Vision Transformer.
+We conduct experiments to verify whether our method can be applied to other architectures, such as Vision Transformer (ViT) (Dosovitskiy et al, 2021).
+Specifically, we adopt ViT-B/16 as the backbone and compare the baseline TSD with our approach.
+The results, shown in Table 5, demonstrate that our method achieves consistent improvements over TSD.
+On the PACS dataset, our method improves the accuracy from 90.20% (TSD) to 91.89%, while on Office-Home, it further boosts performance from 81.80% to 87.32%.
+These gains highlight that our approach is not limited to convolutional networks but can also optimize transformer-based architectures, making it a versatile solution for various backbone choices.
+6 Conclusion
 In this work, we introduce Few-Shot Test Time Adaptation (FS-TTA), a novel setting that diverges from traditional TTA by leveraging the few-shot support set to improve adaptation to the target.
@@
 Looking ahead, we aspire to expand FS-TTA beyond current scope by investigating potential real-world tasks, instead of limiting to image recognition.
+7 Data Availability
+The datasets used in the experiments are all public available.
+Each dataset is under a permissive license that allows usage for research purposes.
+The PACS is available at https://github.com/MachineLearning2020/Homework3-PACS/tree/master/PACS.
+The Office-Home is available at https://www.hemanthdv.org/officeHomeDataset.html.
+The DomainNet is available at http://ai.bu.edu/M3SDA/.

### v4

@@
-
+Enhancing Test Time Adaptation with Few-shot Guidance
 Abstract
-Deep neural networks often experience significant performance degradation when encountering domain shifts between training (source) and test (target) data.
-To mitigate this issue, Test Time Adaptation (TTA) methods have been introduced to adapt a pre-trained source model for handling out-of-distribution streaming target data.
-While these methods provide some improvement, they often lack a dependable mechanism for domain shift correction, which can be unpredictable in real-world applications.
-To address this challenge, we propose the Few-Shot Test Time Adaptation (FS-TTA), a novel and practical setting that builds upon TTA by leveraging a few-shot support set.
-Following the principle of few inputs, big gains, FS-TTA effectively reduces the uncertainty of blind adaptation in previously unseen target domains.
-Furthermore, we introduce a two-stage framework to effectively tackle FS-TTA.
-The first stage involves fine-tuning the pre-trained source model using the few-shot support set while incorporating a feature diversity augmentation module to prevent overfitting.
-The second stage utilizes test time adaptation guided by a prototype memory bank, which facilitates the generation of high-quality pseudo-labels for model adaptation.
-Extensive experiments conducted on three cross-domain classification benchmarks validate the effectiveness, reliability, and superior performance of our proposed FS-TTA and framework.
-keywords:
-Test Time Adaptation, Few-Shot, Domain Shift, Transfer Learning.
+Deep neural networks often suffer severe performance degradation when deployed in target domains that differ from the training data.
+Test-Time Adaptation (TTA) seeks to alleviate this issue by updating a pre-trained model using only streaming unlabeled target data.
+However, the complete absence of supervision makes it difficult for TTA to correct misaligned decision boundaries, often resulting in unstable or even harmful updates under distribution shift.
+We propose Few-Shot Test-Time Adaptation (FS-TTA), a practical extension of TTA that leverages a small labeled support set from the target domain prior to deployment.
+Even minimal supervision substantially reduces the uncertainty of blind adaptation and provides the boundary information needed for more reliable model updates.
+Building on this insight, we introduce a unified two-stage adaptation framework.
+Stage I performs fine-tuning for boundary alignment using the labeled support samples, further enhanced by feature-level diversity to improve robustness in the low-shot regime.
+Stage II conducts test-time distribution refinement using the unlabeled target stream, where reliable pseudo-labels are produced via a prototype memory mechanism to ensure stable online adaptation.
+Extensive experiments on PACS, OfficeHome, and DomainNet demonstrate that the proposed FS-TTA setting and framework deliver consistently superior performance and significantly improve adaptation reliability over state-of-the-art TTA approaches.
+I Introduction
+In recent years, deep neural networks have exhibited remarkable capabilities in representation learning.
+However, their performance relies heavily on the assumption that the distributions of training (source) and test (target) data are identical [long2015learning, ganin2015unsupervised, li2017deeper].
+In real-world deployment, such a distribution shift is inevitable, as it is practically impossible to collect and annotate data for all possible environments in advance of training.
+Besides, this distribution shift can significantly degrade the performance of the deployed source model.
+To address the aforementioned issues, numerous studies have proposed solutions via domain adaptation [long2015learning, tzeng2017adversarial, MDD, xiao2021dynamic, xin2023self] and domain generalization [volpi2018generalizing, zhou2021domain, kim2021selfreg, sicilia2023domain].
+While these approaches have demonstrated impressive performance gains on realistic benchmarks, a considerable gap remains between their problem settings and practical application scenarios.
+Domain adaptation relies on the impractical assumption that target domain data are available and participate in the source training process.
+In contrast, domain generalization aims to directly enhance the generalization of the source model without exploring the target domain data, even if they can be obtained during the test time.
 Figure 1: Test Time Adaptation (TTA) vs.
@@
 FS-TTA incorporates a small number of labeled target samples, which can be easily collected offline before deployment with minimal annotation effort, in addition to the unlabeled target data used in TTA.
-The results for TTA are based on the performance of TENT (Wang et al, 2021) on the OfficeHome (Venkateswara et al, 2017).
-1 Introduction
-In recent years, deep neural networks have exhibited remarkable capabilities in representation learning.
-However, their performance relies heavily on the assumption that the distributions of training (source) and test (target) data are identical (Long et al, 2015; Ganin and Lempitsky, 2015; Li et al, 2017).
-In real-world deployment, such a distribution shift is inevitable, as it is practically impossible to collect and annotate data for all possible environments in advance of training.
-Besides, this distribution shift can significantly degrade the performance of the deployed source model.
-To address the aforementioned issues, numerous studies have proposed solutions via domain adaptation (Long et al, 2015; Tzeng et al, 2017; Zhang et al, 2019; Xiao and Zhang, 2021; Xin et al, 2023) and domain generalization (Volpi et al, 2018; Zhou et al, 2021; Kim et al, 2021; Sicilia et al, 2023).
-While these approaches have demonstrated impressive performance gains on realistic benchmarks, a considerable gap remains between their problem settings and practical application scenarios.
-Domain adaptation relies on the impractical assumption that target domain data are available and participate in the source training process.
-In contrast, domain generalization aims to directly enhance the generalization of the source model without exploring the target domain data, even if they can be obtained during the test time.
-In order to overcome these limitations of domain adaptation/generalization and protect the privacy of the source data, TENT (Wang et al, 2021) introduces fully test time adaptation (TTA).
+The results for TTA are based on the performance of TENT [wang2020tent] on the OfficeHome [venkateswara2017deep].
+In order to overcome these limitations of domain adaptation/generalization and protect the privacy of the source data, TENT [wang2020tent] introduces fully test time adaptation (TTA).
 TTA aims to adapt a pre-trained source model to the target domain using input mini-batch data during the test time, without relying on source data or supervision.
@@
 To serve this purpose, TENT employs test-time entropy minimization to reduce the generalization error on shifted target data.
-Additionally, extensive research has sought to improve TTA through various approaches such as pseudo-labeling (Iwasawa and Matsuo, 2021; Wang et al, 2023), consistency regularization (Boudiaf et al, 2022), and anti-forgetting regularization (Niu et al, 2022).
+Additionally, extensive research has sought to improve TTA through various approaches such as pseudo-labeling [IwasawaM21, wang2023feature], consistency regularization [boudiaf2022parameter], and anti-forgetting regularization [niu2022efficient].
 While these methods can perform model adaptation during the test time, they encounter three primary challenges:
@@
 In scenarios where high-quality, relevant unlabeled data is scarce or not fully representative of the entire target domain, TTA methods may face difficulties in achieving optimal performance, highlighting a major limitation in their application across various real-world settings.
+The fundamental difficulty of fully unsupervised test-time adaptation is that the model must adjust to distribution shifts while blindly exploring the target domain without any reliable supervisory signal.
+This often results in unstable updates, making such methods difficult to deploy in real applications where robustness is essential.
+In many practical systems, it is both feasible and common to obtain a few labeled target samples before deployment, as domain experts or users can annotate representative examples during system setup.
+Even such minimal supervision can provide meaningful guidance during the initial stage of adaptation.
+This naturally leads to the question: If given limited supervisory information from the target domain, could the adaptation performance be improved?
 Figure 2: Performance comparison across different adaptation strategies on OfficeHome.
 One-shot fine-tuning with a single labeled sample per class already surpasses TENT, showing the effectiveness of minimal supervision.
-The fundamental reason for these challenges lies in the blind exploration of the target domain without any supervisory signal under domain shift.
-In practical scenarios, however, it is often feasible to obtain a small number of labeled target samples with minimal annotation cost.
-For instance, a few representative examples can be labeled offline before deployment by domain experts or system users.
-This leads us to ask: If given limited supervisory information from the target domain, could the adaptation performance be improved?
 To answer this question, we test the one-shot situation, as shown in Figure 2.
@@
 We find that the performance is easily improved compared to TENT, which shows that little supervision information can be more effective than a large amount of unsupervised information.
-Based on the aforementioned findings, we introduce the Few-Shot Test Time Adaptation (FS-TTA), encapsulating the concept of few inputs, big gains.
-As illustrated in Figure 1, by integrating a few-shot support set from the target domain prior to model adaptation, FS-TTA effectively reduces domain shift while retaining the source-free and online characteristics inherent to TTA.
-It is noted that such setting is particularly beneficial in scenarios where precision and reliability are paramount, such as medical image analysis and autonomous driving, even spending a few extra labeling costs.
-To solve the FS-TTA, we develop an effective framework.
-For domain shift correction, we first fine-tune the pre-trained source model with the few-shot support set, fostering initial adaptation to the target.
-To prevent overfitting, we propose Feature Diversity Augmentation (FDA) to generate new features.
-During the test time, we employ a self-training strategy, which involves assigning pseudo-labels to unlabeled online mini-batches and using these labels to further update the model online.
-Furthermore, in order to reduce the impact of noisy pseudo-labels on the model, we propose Entropy Filter and Consistency Filter.
-The former filters out high-entropy samples with low confidence, and the latter is achieved through dual-branch prediction consistency.
-The experimental results across various cross-domain image recognition datasets show that our FS-TTA method significantly surpasses the performance of state-of-the-art methods and other baselines.
-To sum up, our main contributions are as follows:
-  • Emerging research direction: We highlight the setting of the Few-Shot Test Time Adaptation (FS-TTA), where a limited few-shot support set is available prior to test-time adaptation.
-Leveraging these few-shot samples enables more effective mitigation of domain shift.
-  • Innovative framework: We propose a meticulously designed framework to address FS-TTA, incorporating fine-tuning of the pre-trained source model with a Feature Diversity Augmentation (FDA) module and performing test-time adaptation through high-quality pseudo-labeled samples in a self-training manner.
+Building on this insight, we propose Few-Shot Test-Time Adaptation (FS-TTA), an adaptation framework that fully exploits the complementary roles of the labeled support set and the unlabeled test-time stream, encapsulating the principle of few inputs, big gains.
+The support set not only alleviates the initial domain shift, but also provides a reliable foundation for subsequent adaptation.
+In Stage I: Fine-tuning for Boundary Alignment, we use the labeled support samples to adjust the source model toward the target domain and reinforce this process with a Feature Diversity Augmentation (FDA) strategy, which regularizes the low-shot update and yields a more reliable initialization.
+On top of this initialization, Stage II: Test-Time Adaptation for Distribution Refinement, leverages the unlabeled test-time stream to further adapt the model.
+The broader distributional information contained in the stream complements the limited support set, while reliable pseudo-labels, selected according to confidence and consistency criteria, ensure that online updates remain stable and resistant to drift under domain shift.
+  • Emerging research direction: We formalize Few-Shot Test-Time Adaptation (FS-TTA), a realistic setting where minimal target supervision is available prior to deployment.
+This setting bridges the gap between fully unsupervised TTA and traditional adaptation paradigms, highlighting the value of incorporating limited but reliable target-domain signals.
+  • Innovative framework: We show that FS-TTA naturally decomposes into a unified framework consisting of boundary-level correction and distribution-level refinement, providing a coherent perspective on how labeled and unlabeled target-domain information should be jointly leveraged for stable and reliable adaptation.
   • State-of-the-art performance: Extensive empirical evaluations on multiple cross-domain classification benchmarks validate the effectiveness of our framework.
 Compared to the current state-of-the-art TTA methods, our approach achieves performance gains of 2.0% on PACS, 7.8% on OfficeHome, and 3.9% on DomainNet.
-Table 1: Comparison with various adaptation settings, where $s$ and $t$ denote source domain and target domain, respectively.
+TABLE I: Comparison with various adaptation settings.
+where $s$ and $t$ denote source domain and target domain.
 $L^{d}$ and $U^{d}$ denote labeled datasets and unlabeled datasets from domain $d$.
 “Online” means that adaptation can predict a batch of incoming test samples immediately.
-“$k$” represents the number of samples per class.
+“kk” represents the number of samples per class.
 “C” indicates the number of classes for the target domain.
@@
 Source domain(s) Target domain Size of available target data
-Domain Generalization $L^{s_{1}},\dots,L^{s_{N}}$ 0
+Domain Generalization $L^{s_{1}},\dots,L^{s_{N}}$ - 0
 Source-Free Domain Adaptation Pre-trained model on $L^{s_{1}},\dots,L^{s_{N}}$ Entire $U^{t}$ $\|U^{t}\|$
 Few-Shot Transfer Learning Pre-trained model on $L^{s_{1}},\dots,L^{s_{N}}$ Few-shot support set $L^{spt}\subset L^{t}$ $k\times\text{C}$
-[…]
-[…]
-2 Related Work
-2.1 Domain Generalization
+Test Time Adaptation Pre-trained model on $L^{s_{1}},\dots,L^{s_{N}}$ Mini-batch $U^{t}$ $|\text{mini-batch}|$, typically 128
+Few-Shot Test Time Adaptation Pre-trained model on $L^{s_{1}},\dots,L^{s_{N}}$ Few-shot support set $L^{spt}\subset L^{t}$ and mini-batch $U^{t}$ $k\times\text{C}$ and $|\text{mini-batch}|$
+II Related Work
+II-A Domain Generalization
 Domain Generalization (DG) aims to train models on multiple related but distinct source domains to ensure effective performance on unseen target domains.
-To enhance robustness, DG techniques often employ strategies such as data augmentation (Huang et al, 2021; Volpi et al, 2018) and data generation (Zhou et al, 2021; Robey et al, 2021) to introduce greater diversity during training.
+To enhance robustness, DG techniques often employ strategies such as data augmentation [huang2021fsdr, volpi2018generalizing] and data generation [zhou2021domain, robey2021model] to introduce greater diversity during training.
 Other prevalent approaches leverage representation learning to extract domain-invariant features.
-This includes kernel-based methods (Li et al, 2018b) that project data into a shared feature space, domain adversarial learning (Sicilia et al, 2023) that aligns distributions via adversarial objectives, and invariant risk minimization (Krueger et al, 2021) which encourages models to perform consistently across domains.
-In addition, self-supervised (Kim et al, 2021) and meta-learning-based techniques (Chen et al, 2022) have been explored to further improve generalization.
+This includes kernel-based methods [li2018domain] that project data into a shared feature space, domain adversarial learning [sicilia2023domain] that aligns distributions via adversarial objectives, and invariant risk minimization [krueger2021out] which encourages models to perform consistently across domains.
+In addition, self-supervised [kim2021selfreg] and meta-learning-based techniques [chen2022discriminative] have been explored to further improve generalization.
 However, without exposure to the target domain, generalization remains inherently limited.
-2.2 Source-Free Domain Adaptation
+II-B Source-Free Domain Adaptation
 Source-Free Domain Adaptation (SFDA) aims to adapt a pre-trained source model to an unlabeled target domain while ensuring that no source data is accessed during the adaptation process.
@@
 SFDA techniques can be broadly categorized into two main approaches: pseudo-label strategies and generative methods.
-The former leverages target pseudo-labels to facilitate self-training, thereby enabling implicit adaptation without requiring explicit supervision (Tanwisuth et al, 2021; Ahmed et al, 2021; Liang et al, 2021; Xin et al, 2023).
-The latter employs generative models to synthesize target-style training data, allowing the model to bridge the domain gap through data augmentation and distribution alignment (Qiu et al, 2021; Liu et al, 2021).
+The former leverages target pseudo-labels to facilitate self-training, thereby enabling implicit adaptation without requiring explicit supervision [tanwisuth2021prototype, ahmed2021unsupervised, liang2021source, xin2023self].
+The latter employs generative models to synthesize target-style training data, allowing the model to bridge the domain gap through data augmentation and distribution alignment [qiu2021source, liu2021source].
 Similar to SFDA, our proposed Few-Shot Test Time Adaptation (FS-TTA) also maintains the source-free property, ensuring that adaptation is performed without relying on source data while leveraging a small support set to enhance adaptation efficiency.
-2.3 Few-Shot Transfer Learning
+II-C Few-Shot Transfer Learning
 Test Time Adaptation (TTA) aims to adapt a pre-trained source model on-the-fly during inference to mitigate distribution shifts.
-Early TTA methods apply self-supervised learning objectives (Sun et al, 2020), but typically require access to training data or modification of the training process.
-TENT (Wang et al, 2021) addresses this by proposing fully test-time adaptation, relying solely on target samples and adapting batch normalization parameters via entropy minimization.
-Subsequent approaches such as (Schneider et al, 2020; Nado et al, 2020) update statistics on each incoming mini-batch, while methods like LAME (Boudiaf et al, 2022) and EATA (Niu et al, 2022) tackle catastrophic forgetting during continual adaptation.
-TSD (Wang et al, 2023) further integrates self-training to selectively update using confident predictions.
+Early TTA methods apply self-supervised learning objectives [sun2020test], but typically require access to training data or modification of the training process.
+TENT [wang2020tent] addresses this by proposing fully test-time adaptation, relying solely on target samples and adapting batch normalization parameters via entropy minimization.
+Subsequent approaches such as [schneider2020improving, nado2020evaluating] update statistics on each incoming mini-batch, while methods like LAME [boudiaf2022parameter] and EATA [niu2022efficient] tackle catastrophic forgetting during continual adaptation.
+TSD [wang2023feature] further integrates self-training to selectively update using confident predictions.
 As a result, adaptation often relies heavily on the quality of incoming test samples.
-2.4 Test Time Adaptation
+II-D Test Time Adaptation
 Test Time Adaptation (TTA) aims to adapt a pre-trained source model during inference to mitigate distribution shifts between training and test domains.
-Early TTA methods address this challenge through self-supervised auxiliary tasks (Sun et al, 2020), which, while effective, often require access to training data or modifications to the training procedure.
-To overcome this limitation, TENT (Wang et al, 2021) proposes fully test-time adaptation by leveraging only target data, updating batch normalization parameters via entropy minimization.
-Building on this, subsequent works (Schneider et al, 2020; Nado et al, 2020) estimate batch normalization statistics dynamically from incoming test batches.
-Other approaches, such as LAME (Boudiaf et al, 2022) and EATA (Niu et al, 2022), focus on preventing catastrophic forgetting during continuous model updates.
-More recently, TSD (Wang et al, 2023) incorporates self-training by selecting confident test samples to guide adaptation.
+Early TTA methods address this challenge through self-supervised auxiliary tasks [sun2020test], which, while effective, often require access to training data or modifications to the training procedure.
+To overcome this limitation, TENT [wang2020tent] proposes fully test-time adaptation by leveraging only target data, updating batch normalization parameters via entropy minimization.
+Building on this, subsequent works [schneider2020improving, nado2020evaluating] estimate batch normalization statistics dynamically from incoming test batches.
+Other approaches, such as LAME [boudiaf2022parameter] and EATA [niu2022efficient], focus on preventing catastrophic forgetting during continuous model updates.
+More recently, TSD [wang2023feature] incorporates self-training by selecting confident test samples to guide adaptation.
 Despite their progress, these methods heavily rely on the quality and stability of online target data.
-2.5 Comparisons with Other Settings
-We compare Few-Shot Test Time Adaptation (FS-TTA) with similar problem settings (details are in the appendix), as illustrated in Table 1.
+II-E Comparisons with Other Settings
+We compare Few-Shot Test Time Adaptation (FS-TTA) with similar problem settings (details are in the appendix), as illustrated in Table I.
   • Compared with Domain Generalization, FS-TTA eliminates the necessity of accessing source data, thereby ensuring the preservation of source data privacy.
@@
 Additionally, FS-TTA demonstrates superior performance in handling challenging scenarios where there are substantial domain shifts, making it a more robust and reliable solution.
-3 Preliminary
-3.1 Instance Normalization
-Instance Normalization (Ulyanov et al, 2016) is a normalization technique widely used in deep neural network architectures, especially in the context of style transfer and generative models.
+Figure 3: Illustration of our two-stage framework.
+In Stage I, we employ the few-shot support set to fine-tune the source model.
+To prevent overfitting, we propose FDA module.
+In Stage II, we maintain a prototype memory bank to guide test time adaptation.
+In order to update the prototype memory bank and model with effective samples, we propose the entropy filter and consistency selection modules.
+III Preliminary
+III-A Instance Normalization
+Instance Normalization [ulyanov2016instance] is a normalization technique widely used in deep neural network architectures, especially in the context of style transfer and generative models.
 Let us consider a batch images with size $N\times C\times H\times W$, where $N$ is the batch size, $C$ is the number of channels, and $H$ and $W$ are the height and width of the images.
@@
 where $\gamma,\beta\in R^{C}$ are learnable transformation parameters.
-3.2 Class Prototype
+III-B Class Prototype
 The class prototype is a representative point in the feature space that summarizes the key characteristics of a class.
@@
 where $P_{c}\in R^{d}$. The class prototype plays an important role in few-shot scenarios.
-Figure 3: Illustration of our two-stage framework.
-In Stage I, we employ the few-shot support set to fine-tune the source model.
-To prevent overfitting, we propose FDA module.
-In Stage II, we maintain a prototype memory bank to guide test time adaptation.
-In order to update the prototype memory bank and model with effective samples, we propose the entropy filter and consistency selection modules.
-4 Method
-4.1 Problem Setting
+IV Method
+IV-A Problem Setting
 Considering a typical scenario where a source model $f_{\theta_{s}}$ is equipped with parameters $\theta_{s}$ and trained on source datasets $\mathcal{D}{s_{1}},\mathcal{D}{s_{2}},…,\mathcal{D}{s_{n}}$, our objective is to adapt this pre-trained model to a target domain $D_{t}$ without accessing source data.
@@
 Notably, the support set $S$ can be acquired offline prior to deployment, and in many real-world applications, collecting such limited supervision is both feasible and cost-efficient.
-4.2 Stage I: Model Adaptation via Fine-Tuning
-To significantly and swiftly enhance the initialization performance of the pre-trained source model in the target domain and minimize domain shifts, we design to fine-tune the pre-trained source model using the few-shot support set.
+IV-B Stage I: Boundary Alignment via Few-Shot Fine-Tuning
+A major challenge in test-time adaptation is that the source model’s decision boundary does not align with the target-domain distribution.
+The few-shot support set provides the only unbiased supervision capable of correcting this boundary misalignment.Therefore, Stage I aims to align the model boundary with the target domain by fine-tuning the pre-trained source model on these labeled samples.
 Given the limited number of samples per class, there is a potential risk of overfitting during the fine-tuning process.
@@
 Feature Diversity Augmentation (FDA).
-Prior research (Zhou et al, 2021) has demonstrated a significant association between feature statistics and image style, which is intricately linked to data distribution within the field of computer vision.
+Prior research [zhou2021domain] has demonstrated a significant association between feature statistics and image style, which is intricately linked to data distribution within the field of computer vision.
 To increase style diversity while preserving semantic consistency, we introduce Feature Diversity Augmentation (FDA), a feature-level data augmentation technique that simulates various image styles without altering the original class labels.
@@
 Secondly, FDA generates the mixtures of feature statistics:
-$\bm{\gamma}_{\text{mix }}=\lambda\sigma_{i}+(1-\lambda)\sigma_{j},$ (5)
-$\bm{\beta}_{\text{mix }}=\lambda\mu_{i}+(1-\lambda)\mu_{j}.$ (6)
+$\boldsymbol{\gamma}_{\text{mix }}=\lambda\sigma_{i}+(1-\lambda)\sigma_{j},$ (5)
+$\boldsymbol{\beta}_{\text{mix }}=\lambda\mu_{i}+(1-\lambda)\mu_{j}.$ (6)
 In this case, $\lambda$ denotes the mixing ratio coefficient. Ultimately, the mixtures of feature statistics are applied to the feature map $f_{i}$ via instance normalization:
-$f_{i}^{\prime}=\bm{\gamma}_{\operatorname{mix}}\odot\frac{f_{i}-\mu_{i}}{\sigma_{i}}+\bm{\beta}_{\operatorname{mix}},$ (7)
+$f_{i}^{\prime}=\boldsymbol{\gamma}_{\operatorname{mix}}\odot\frac{f_{i}-\mu_{i}}{\sigma_{i}}+\boldsymbol{\beta}_{\operatorname{mix}},$ (7)
 where $f_{i}^{\prime}$ represents the newly generated feature map.
@@
 Specifically, the few-shot support set is processed through $f_{\theta_{s}}$ to minimize a supervised loss, defined as:
-$\mathcal{L}_{\text{cls}}=-\sum_{i=1}^{k*C}\mathcal{H}\left(y_{i},p\left(\hat{y_{i}}\mid\bm{s}_{i}\right)\right),$ (8)
+$\mathcal{L}_{\text{cls}}=-\sum_{i=1}^{k*C}\mathcal{H}\left(y_{i},p\left(\hat{y_{i}}\mid\boldsymbol{s}_{i}\right)\right),$ (8)
 where $\mathcal{H}(\cdot)$ is the cross-entropy loss.
 The term $y_{i}$ is the ground-truth label of $s_{i}$, indicating one of sample from few-shot support set, and $C$ represents categories of the target.
-4.3 Stage II: Test Time Adaptation
-During this stage, a mini-batch of unlabeled samples, denoted as $x=\{x_{1},x_{2},..,x_{B}\}$, online arrives.
+IV-C Stage II: Test Time Adaptation
+After obtaining a target-aware initialization from Stage I, the model encounters a stream of unlabeled target samples $x=\{x_{1},x_{2},\ldots,x_{B}\}$ during deployment.
+These samples provide additional distributional insights beyond what the support set can offer.
+Stage II aims to refine the model online by leveraging this unlabeled stream while ensuring stable updates under distribution shift.
 The central concept of this stage is to employ a self-training strategy to update the fine-tuned source model online, enabling it to fully adapt to the target domain.
@@
 To address this issue, we propose two modules to produce high quality pseudo-labels.
-The first is entropy filter, which screens out unreliable samples using Shannon entropy (Shannon, 2001).
+The first is entropy filter, which screens out unreliable samples using Shannon entropy [shannon2001mathematical].
 Typically, samples with higher entropy are considered to have lower prediction confidence.
@@
 To dynamically update the model using online mini-batch target, it is crucial to filter out noisy samples, as they may be assigned to incorrect classes, resulting in inaccurate prototype computation.
-In this regard, we propose the Entropy Filter, which employs Shannon entropy (Shannon, 2001) to select confident samples in the mini-batch.
+In this regard, we propose the Entropy Filter, which employs Shannon entropy [shannon2001mathematical] to select confident samples in the mini-batch.
 For an sample $x_{i}$, its entropy can be computed as:
 $H(p_{i})=-\sum(p_{i})\cdot log(p_{i}).$ (9)
-Based on the insights from previous work (Wang et al, 2021), high entropy samples should be filtered out, as lower entropy typically indicates higher accuracy.
+Based on the insights from previous work [wang2020tent], high entropy samples should be filtered out, as lower entropy typically indicates higher accuracy.
 Consequently, we sort the entropy of all samples in the mini-batch and select the top $\alpha\%$ samples with lower entropy, donated as $\hat{x}=\{\hat{x}_{1},\hat{x}_{2},\ldots,\hat{x}_{\lfloor\alpha\cdot B\rfloor}\}$.
@@
 where $m_{c_{t}}$ represents the $c$-th class prototype at time $t$, and $\beta$ represents the sliding update coefficient.
-Table 2: Table 2: Comparison with test-time adaptation methods on three datasets with ResNet-50 backbone.
-FS-TTA achieves consistent improvements over TSD (Wang et al, 2021), the strongest baseline method.
-OfficeHome PACS DomainNet
-Method Art Clip Prod Real Avg. Art Cart Phot Sket Avg. Avg.
-Test time adaptation methods
-ERM (Vapnik, 1999) 60.7 55.7 76.2 76.8 67.4 82.5 80.8 94.0 80.9 84.5 45.2
-BN (Nado et al, 2020) 58.2 55.6 75.1 75.5 66.1 83.2 84.9 94.0 77.9 85.0 43.3
-TENT (Wang et al, 2021) 60.6 58.7 76.5 76.1 68.0 85.2 86.7 94.9 82.9 87.4 44.7
-T3A (Iwasawa and Matsuo, 2021) 61.2 56.7 78.0 77.3 68.3 84.0 82.3 95.0 82.7 86.0 46.1
-ETA (Niu et al, 2022) 58.4 55.8 75.2 75.5 66.2 83.2 84.9 94.0 77.9 85.0 46.1
-LAME (Boudiaf et al, 2022) 58.7 55.6 75.1 75.4 66.2 84.9 85.5 95.0 80.9 86.6 43.2
-TSD (Wang et al, 2023) 62.3 57.5 77.5 77.5 68.7 87.6 88.7 96.1 85.0 89.4 47.7
-PROGRAM (Sun et al, 2024) 63.4 54.3 77.2 77.2 68.0 87.2 84.1 96.9 76.4 86.2 43.3
-DEYO (Lee et al, 2024) 63.8 54.9 76.4 77.3 68.1 88.4 85.2 97.1 82.3 88.2 42.5
-Fine-tuning + Test time adaptation methods
-FT+TENT (Wang et al, 2021) 68.8 65.5 79.8 78.5 73.2 87.0 86.9 95.2 83.6 88.2 45.4
-FT+TSD (Wang et al, 2023) 70.5 65.1 80.3 79.2 73.8 88.3 88.6 96.5 85.9 89.8 48.5
-FS-TTA 73.2 68.3 83.0 81.6 76.5 90.4 89.7 97.6 87.8 91.4 51.6
-$\Delta_{up}$ over TSD (+10.9$\uparrow$ (+10.8$\uparrow$ (+5.5$\uparrow$ (+4.1$\uparrow$ (+7.8$\uparrow$ (+2.8$\uparrow$ (+1.0$\uparrow$ (+1.5$\uparrow$ (+2.8$\uparrow$ (+2.0$\uparrow$ (+3.9$\uparrow$
 Test Time Adaptation.
@@
 This strategy can be implemented through a filter mask for samples $x_{j}$ as follows:
-$\mathcal{M}_{j}=\mathbbm{1}[\arg\max p_{j}=\arg\max\hat{p}_{j}].$ (13)
+$\mathcal{M}_{j}=\mathds{1}[\arg\max p_{j}=\arg\max\hat{p}_{j}].$ (13)
 Ultimately, we can update the model using reliable samples, and the loss can be formulated as follows:
@@
 It’s noteworthy that our self-training process does not involve specifying any threshold, which enhances the model’s generalizability.
-5 Experiment
-5.1 Experimental Settings
+TABLE II: Comparison with test-time adaptation methods on three datasets with ResNet-50 backbone. FS-TTA achieves consistent improvements over TSD [wang2020tent], the strongest baseline method.
+OfficeHome PACS DomainNet
+Method Art Clip Prod Real Avg. Art Cart Phot Sket Avg. Avg.
+Test time adaptation methods
+ERM [vapnik1999overview] 60.7 55.7 76.2 76.8 67.4 82.5 80.8 94.0 80.9 84.5 45.2
+BN [nado2020evaluating] 58.2 55.6 75.1 75.5 66.1 83.2 84.9 94.0 77.9 85.0 43.3
+TENT [wang2020tent] 60.6 58.7 76.5 76.1 68.0 85.2 86.7 94.9 82.9 87.4 44.7
+T3A [IwasawaM21] 61.2 56.7 78.0 77.3 68.3 84.0 82.3 95.0 82.7 86.0 46.1
+ETA [niu2022efficient] 58.4 55.8 75.2 75.5 66.2 83.2 84.9 94.0 77.9 85.0 46.1
+LAME [boudiaf2022parameter] 58.7 55.6 75.1 75.4 66.2 84.9 85.5 95.0 80.9 86.6 43.2
+PROGRAM [sun2024program] 63.4 54.3 77.2 77.2 68.0 87.2 84.1 96.9 76.4 86.2 43.3
+DEYO [lee2024entropy] 63.8 54.9 76.4 77.3 68.1 88.4 85.2 97.1 82.3 88.2 42.5
+TSD [wang2023feature] 62.3 57.5 77.5 77.5 68.7 87.6 88.7 96.1 85.0 89.4 47.7
+Fine-tuning + Test time adaptation methods
+FT+TENT [wang2020tent] 68.8 65.5 79.8 78.5 73.2 87.0 86.9 95.2 83.6 88.2 45.4
+FT+TSD [wang2023feature] 70.5 65.1 80.3 79.2 73.8 88.3 88.6 96.5 85.9 89.8 48.5
+FS-TTA 73.2 68.3 83.0 81.6 76.5 90.4 89.7 97.6 87.8 91.4 51.6
+[…]
+V Experiment
+V-A Experimental Settings
 Dataset. To evaluate the effectiveness of our proposed setting and method, we conduct experiments on three cross-domain benchmarks.
-  • PACS (Li et al, 2017) consists of 9,991 images spanning four distinct domains: Art, Cartoon (Cart), Photo (Phot), and Sketch (Sket).
+  • PACS [li2017deeper] consists of 9,991 images spanning four distinct domains: Art, Cartoon (Cart), Photo (Phot), and Sketch (Sket).
 Each domain contains seven object categories: dog, elephant, giraffe, guitar, horse, house, and person.
-  • Office-Home (Venkateswara et al, 2017) comprises 15,588 images distributed across four domains: Art, Clipart (Clip), Product (Prod), and Real-World (Real), with each domain encompassing 65 image categories.
-  • DomainNet (Peng et al, 2019) is a large-scale dataset containing six domains: Clipart (Clip), Infograph (Info), Painting (Pain), Quickdraw (Quic), Real, Sketch (Sket), comprising a total of 586,575 images across 345 classes.
+  • Office-Home [venkateswara2017deep] comprises 15,588 images distributed across four domains: Art, Clipart (Clip), Product (Prod), and Real-World (Real), with each domain encompassing 65 image categories.
+  • DomainNet [peng2019moment] is a large-scale dataset containing six domains: Clipart (Clip), Infograph (Info), Painting (Pain), Quickdraw (Quic), Real, Sketch (Sket), comprising a total of 586,575 images across 345 classes.
 Implementation Details.
-In our main experiments, we employ ResNet-50(He et al, 2016), pre-trained on ImageNet-1k(Russakovsky et al, 2015), as the backbone model, as it is widely adopted in the test-time adaptation literature.
-For source model training, we follow the leave-one-domain-out protocol, as recommended by prior studies (Wang et al, 2023; Zhou et al, 2021), treating one domain as the unlabeled target and the rest as source domains.We set the batch size to 32 for each source domain and use a learning rate of 5e-5.
+In our main experiments, we employ ResNet-50[he2016deep], pre-trained on ImageNet-1k[russakovsky2015imagenet], as the backbone model, as it is widely adopted in the test-time adaptation literature.
+For source model training, we follow the leave-one-domain-out protocol, as recommended by prior studies [wang2023feature, zhou2021domain], treating one domain as the unlabeled target and the rest as source domains.We set the batch size to 32 for each source domain and use a learning rate of 5e-5.
 Both the dropout probability and weight decay are set to zero.
-The source model is trained for 5,000 iterations, except for DomainNet, where we extend training to 15,000 iterations, following the methodology in (Cha et al, 2021).
+The source model is trained for 5,000 iterations, except for DomainNet, where we extend training to 15,000 iterations, following the methodology in [swad].
 All images are resized to 224 × 224, and data augmentation is applied during source domain training, including random cropping, horizontal flipping, color jittering, and intensity adjustments.
@@
 We carry out all experiments on NVIDIA V100 GPUs.
-Figure 4: Comprehensive comparison between our method and the state-of-the-art method in DG/TTA settings on DomainNet.
-Table 3: Compared with existing DG and SFDA methods on OfficeHome and DomainNet.
+TABLE III: Compared with existing DG and SFDA methods on OfficeHome and DomainNet.
 OfficeHome DomainNet
@@
 Domain generalization methods
-ERM (Vapnik, 1999) 60.7 55.7 76.2 76.8 67.4 64.8 22.1 51.8 13.8 64.7 54.0 45.2
-DNA (Chu et al, 2022) 67.7 57.7 78.9 80.5 71.2 66.1 23.0 54.6 16.7 65.8 56.8 47.2
-PCL (Yao et al, 2022) 67.3 59.9 78.7 80.7 71.6 67.9 24.3 55.3 15.7 66.6 56.4 47.7
-SWAD (Cha et al, 2021) 66.1 57.7 78.4 80.2 70.6 66.1 22.4 53.6 16.3 65.5 56.2 46.7
+ERM [vapnik1999overview] 60.7 55.7 76.2 76.8 67.4 64.8 22.1 51.8 13.8 64.7 54.0 45.2
+DNA [chu2022dna] 67.7 57.7 78.9 80.5 71.2 66.1 23.0 54.6 16.7 65.8 56.8 47.2
+PCL [yao2022pcl] 67.3 59.9 78.7 80.7 71.6 67.9 24.3 55.3 15.7 66.6 56.4 47.7
+SWAD [swad] 66.1 57.7 78.4 80.2 70.6 66.1 22.4 53.6 16.3 65.5 56.2 46.7
 Source-free domain adaptation methods
-F-mix (Kundu et al, 2022) 72.6 67.4 85.9 83.6 77.4 75.4 24.6 57.8 23.6 65.8 58.5 51.0
+F-mix [KunduKBMKJR22] 72.6 67.4 85.9 83.6 77.4 75.4 24.6 57.8 23.6 65.8 58.5 51.0
 FS-TTA 73.2 68.3 83.0 81.6 76.5 68.6 30.8 56.4 24.2 69.1 60.2 51.6
-SWAD + FS-TTA 77.4 71.1 86.4 84.2 79.8 - - - - - - -
+SWAD + FS-TTA 77.4 71.1 86.4 84.2 79.8 69.2 31.5 57.9 25.1 70.8 62.0 52.8
 Baselines.
-We compare our method with various test-time adaptation (TTA) approaches, including BN(Nado et al, 2020), TENT(Wang et al, 2021), ETA(Niu et al, 2022), T3A(Iwasawa and Matsuo, 2021), LAME(Boudiaf et al, 2022), TSD(Wang et al, 2023),PROGRAM (Sun et al, 2024) and DEYO (Lee et al, 2024).
+We compare our method with various test-time adaptation (TTA) approaches, including BN[nado2020evaluating], TENT[wang2020tent], ETA[niu2022efficient], T3A[IwasawaM21], LAME[boudiaf2022parameter], TSD[wang2023feature],PROGRAM [sun2024program] and DEYO [lee2024entropy].
 Additionally, we establish new baselines by integrating fine-tuning with existing TTA methods to ensure a more comprehensive comparison.
-Furthermore, we compare our approach with selected methods from domain generalization, source-free domain adaptation, including DNA(Chu et al, 2022), PCL(Yao et al, 2022), SWAD(Cha et al, 2021), and F-mix(Kundu et al, 2022).
-Finally, we set up a comparison with the few-shot transfer learning methods, including AdaBN (Li et al, 2016), $L^{2}$ (Li et al, 2018a), DELTA (Li et al, 2019), FLUTE (Triantafillou et al, 2021), LCCS (Zhang et al, 2022).
+Furthermore, we compare our approach with selected methods from domain generalization, source-free domain adaptation, including DNA[chu2022dna], PCL[yao2022pcl], SWAD[swad], and F-mix[KunduKBMKJR22].
+Finally, we set up a comparison with the few-shot transfer learning methods, including AdaBN [li2016AdaBN], $L^{2}$ [li2018explicitTL], DELTA [Li2019DELTADL], FLUTE [triantafillou2021flute], LCCS [zhang2022few].
 For a global overview, we compare our method with state-of-the-art method in various settings, as shown in Figure 4.
-5.2 Performance Comparisons
+V-B Performance Comparisons
 Comparison with TTA methods.
-Table 2 details the comparison results between our method and various TTA methods on the Office-Home and PACS datasets, as well as the final results of DomainNet (detailed in Table 3).
+Table II details the comparison results between our method and various TTA methods on the Office-Home and PACS datasets, as well as the final results of DomainNet (detailed in Table III).
 We observe that our method achieves state-of-the-art performance.
+Figure 4: Comprehensive comparison between our method and the state-of-the-art method in DG/TTA settings on DomainNet.
 Primarily, our approach exhibits a significant enhancement in performance compared to the source model (ERM).
@@
 On the other two datasets, we observe average performance increments of 6.9% (PACS) and 6.4% (DomainNet).
-Table 4: Compared with few-shot transfer learning methods on PACS dataset.
+TABLE IV: Compared with few-shot transfer learning methods on PACS dataset.
 PACS
@@
 Few-shot transfer learning methods
-AdaBN (Li et al, 2016) 85.0 83.5 96.0 78.7 85.8
-$L^{2}$ (Li et al, 2018a) 85.6 84.1 96.4 76.3 85.6
-DELTA (Li et al, 2019) 85.6 83.8 96.5 76.3 85.6
-FLUTE (Triantafillou et al, 2021) 87.2 86.1 97.2 81.7 88.1
-LCCS (Zhang et al, 2022) 87.7 86.9 97.5 83.0 88.8
+AdaBN [li2016AdaBN] 85.0 83.5 96.0 78.7 85.8
+$L^{2}$ [li2018explicitTL] 85.6 84.1 96.4 76.3 85.6
+DELTA [Li2019DELTADL] 85.6 83.8 96.5 76.3 85.6
+FLUTE [triantafillou2021flute] 87.2 86.1 97.2 81.7 88.1
+LCCS [zhang2022few] 87.7 86.9 97.5 83.0 88.8
 FS-TTA 90.4 89.8 97.6 87.9 91.4
@@
 Finally, for a more comprehensive comparison with TTA methods, we construct new baselines by combining fine-tuning with representative TTA approaches.
-Specifically, we select TENT (Wang et al, 2021), as a widely adopted and foundational method in test-time adaptation, and TSD (Wang et al, 2023), which demonstrates state-of-the-art performance across benchmarks.
-According to the results in Table 2, our method achieves an average improvement of 4.2% over Fine-Tuning+TENT and 2.4% over Fine-Tuning+TSD across the three datasets.
+Specifically, we select TENT [wang2020tent], as a widely adopted and foundational method in test-time adaptation, and TSD [wang2023feature], which demonstrates state-of-the-art performance across benchmarks.
+According to the results in Table II, our method achieves an average improvement of 4.2% over Fine-Tuning+TENT and 2.4% over Fine-Tuning+TSD across the three datasets.
 These results highlight the superiority of our framework in migrating to the few-shot TTA setting, benefiting from the proposed FDA module and the support-set-based prototype initialization.
-Figure 5: Ablation Study on (a) effectiveness analysis about two-stage framework and (b) effectiveness analysis about FDA module.
-Figure 6: Comprehensive Sensitivity Analysis of (1) Different Entropy Filter Proportion $\alpha$ and K-shot Number on Model Performance.
+Figure 5: (a) Effectiveness of the two-stage framework.
+Stage I already yields notable gains over the source model, while adding Stage II brings further improvements, demonstrating the complementary roles of boundary alignment and distribution refinement.
+(b) Effectiveness of the FDA module.
+FDA improves few-shot fine-tuning over both the baseline and Mix-up, showing that enhanced feature diversity benefits low-shot adaptation.
 Comparison with DG/SFDA methods.
@@
 To answer this question, we compare our method with several methods in DG and SFDA.
-The results of Office-Home dataset are shown in Table 3.
+The results of Office-Home dataset are shown in Table III.
 It can be seen that our method outperforms the state-of-the-art methods in DG, such as SWAD and PCL.
-Furthermore, equipped with SWAD (SWAD+FS-TTA), our method achieves 79.8% accuracy.
-This result benefits from our adaptation of the model during the test time.
-In comparison to advanced SFDA methods, FS-TTA still achieves satisfactory results.
+In addition, we combine FS-TTA with a SWAD-pretrained model to examine whether our framework can further benefit from a stronger DG-trained initialization.
+The resulting SWAD+FS-TTA model achieves 79.8% accuracy on OfficeHome, indicating that FS-TTA can effectively leverage the flat-minima representations learned by SWAD and yield additional improvements during test-time adaptation.
+This demonstrates that FS-TTA is highly compatible with diverse source-model training paradigms, serving as a plug-and-play adaptation module that can be seamlessly integrated with various DG optimizers to further enhance robustness under domain shift.In comparison to advanced SFDA methods, FS-TTA still achieves satisfactory results.
 It is worth noting that FS-TTA is more flexible in real-world scenarios than SFDA since it adapts the target data in an offline manner, requiring more training loops and resources.
-The results of DomainNet are shown in Table 3.
+The results of DomainNet are shown in Table III.
 The overall performance of FS-TTA outperforms the SFDA methods, suggesting that FS-TTA is more adept at handling challenging tasks.
@@
 To ensure a comprehensive evaluation, we compare our approach with existing few-shot transfer learning methods.
-The results on the PACS dataset are presented in Table 4.
+The results on the PACS dataset are presented in Table IV.
 According to the results, FS-TTA consistently outperforms all baseline methods across different domains, achieving the highest average accuracy of 91.4%, which surpasses the best-performing baseline LCCS (88.8%).
 This result highlights the effectiveness of our approach in adapting to domain shifts and improving classification performance in the few-shot setting.
-5.3 Ablation Study
+V-C Ablation Study
 Effectiveness of two-stage framework.
@@
 As a result, our two-stage framework establishes itself as a robust foundation for the Few-Shot Test Time Adaptation setting, demonstrating its considerable potential in enabling online model adaptation in real-world situations where labeled data is scarce.
-Figure 7: The t-SNE feature visualization of (a) ERM, (b) TENT, (c) LAME, and (d) FS-TTA (Ours).
-Figure 8: Analyzing the Impact of Batch Size on Accuracy and Running Efficiency.
+Figure 6: (a) Sensitivity of pseudo-label filtering to different entropy-filter proportions $\alpha$.
+(b) Effect of K-shot supervision strength on adaptation accuracy.
+(c) Influence of batch size on accuracy during online refinement.
+(d) Influence of batch size on running efficiency.
+Figure 7: t-SNE visualization of target-domain features under different adaptation methods.
+The plots compare the feature distributions produced by (a) ERM, (b) TENT, (c) LAME, and (d) our FS-TTA framework, illustrating how various strategies influence the target-domain representation structure.
 Effectiveness of FDA module.
@@
 To evaluate the impact of $\alpha$, we conduct an experimental analysis on the Office-Home dataset by assigning $\alpha$ to 0, 0.3, 0.6, and 1, respectively.
-The results, as shown in Figure 6(b), demonstrate that $\alpha>0$ yields performance improvements compared to $\alpha=0$ (the source model), highlighting the effectiveness of our proposed framework.
+The results, as shown in Figure 6(a), demonstrate that $\alpha>0$ yields performance improvements compared to $\alpha=0$ (the source model), highlighting the effectiveness of our proposed framework.
 Furthermore, $\alpha=0.3$ and $\alpha=0.6$ perform better than $\alpha=1$ (no filter), indicating the effectiveness of our entropy filter strategy.
@@
 To elucidate the impact of the number of k-shots on our method, we carry out additional ablation experiments within the Office-Home dataset.
-The findings, illustrated in Figure 6, indicate a significant performance enhancement when the shot size ranges from 1 to 10, demonstrating a rapid performance ascension in this few-shot regime.
+The findings, illustrated in Figure 6(b), indicate a significant performance enhancement when the shot size ranges from 1 to 10, demonstrating a rapid performance ascension in this few-shot regime.
 Remarkably, even minimal shot sizes such as 1-shot and 3-shot exhibit substantial effectiveness.
 For instance, the 3-shot configuration achieves a 3.8% performance improvement over the TSD.
+Efficiency analysis.
+In our main experiments, we opt for a mini-batch size of 64.
+To examine the variations in performance and computational efficiency with different batch size during test-time adaptation, we conduct a series of analytical experiments.
+As shown in Figure 6(c), we observe that accuracy experiences a gradual increase as the batch size incrementally grows, reaching a plateau around a batch size of 64.
+In contrast, as shown in Figure 6(d), running time exhibits a decreasing trend as the batch size grows.
+However, beyond a batch size of 64, the running time appears to stabilize.
+Consequently, for real-world applications aiming to achieve a trade-off between accuracy and computational efficiency, we suggest a batch size in the vicinity of 64.
+TABLE V: Comparison of our method with the baseline TSD on both ResNet and ViT-B/16 backbones across the PACS and Office-Home datasets.
+Backbones   PACS Office-Home
+ResNet 84.59 67.37
++ TSD [wang2023feature] 89.41 68.67
++ Ours 91.42 76.49
+ViT-B/16 87.13 79.06
++ TSD [wang2023feature] 90.20 81.80
++ Ours 91.89 87.32
 Qualitative analysis by t-SNE visualization.
@@
 In contrast, our method produces more uniform and aligned feature distribution after adapting to the target domain, as shown in Figure 7(d).
-Efficiency analysis.
-In our main experiments, we opt for a mini-batch size of 64.
-To examine the variations in performance and computational efficiency with different batch size during test-time adaptation, we conduct a series of analytical experiments.
-As shown in Figure 8(a), we observe that accuracy experiences a gradual increase as the batch size incrementally grows, reaching a plateau around a batch size of 64.
-In contrast, as shown in Figure 8(b), running time exhibits a decreasing trend as the batch size grows.
-However, beyond a batch size of 64, the running time appears to stabilize.
-Consequently, for real-world applications aiming to achieve a trade-off between accuracy and computational efficiency, we suggest a batch size in the vicinity of 64.
-Table 5: Comparison of our method with the baseline TSD on both ResNet and ViT-B/16 backbones across the PACS and Office-Home datasets.
-Backbones PACS Office-Home
-ResNet 84.59 67.37
-+ TSD (Wang et al, 2023) 89.41 68.67
-+ Ours 91.42 76.49
-ViT-B/16 87.13 79.06
-+ TSD (Wang et al, 2023) 90.20 81.80
-+ Ours 91.89 87.32
 Scalability on Vision Transformer.
-We conduct experiments to verify whether our method can be applied to other architectures, such as Vision Transformer (ViT) (Dosovitskiy et al, 2021).
+We conduct experiments to verify whether our method can be applied to other architectures, such as Vision Transformer (ViT) [dosovitskiy2020image].
 Specifically, we adopt ViT-B/16 as the backbone and compare the baseline TSD with our approach.
-The results, shown in Table 5, demonstrate that our method achieves consistent improvements over TSD.
+The results, shown in Table V, demonstrate that our method achieves consistent improvements over TSD.
 On the PACS dataset, our method improves the accuracy from 90.20% (TSD) to 91.89%, while on Office-Home, it further boosts performance from 81.80% to 87.32%.
 These gains highlight that our approach is not limited to convolutional networks but can also optimize transformer-based architectures, making it a versatile solution for various backbone choices.
-6 Conclusion
+VI Conclusion
 In this work, we introduce Few-Shot Test Time Adaptation (FS-TTA), a novel setting that diverges from traditional TTA by leveraging the few-shot support set to improve adaptation to the target.
@@
 Looking ahead, we aspire to expand FS-TTA beyond current scope by investigating potential real-world tasks, instead of limiting to image recognition.
-7 Data Availability
-The datasets used in the experiments are all public available.
-Each dataset is under a permissive license that allows usage for research purposes.
-The PACS is available at https://github.com/MachineLearning2020/Homework3-PACS/tree/master/PACS.
-The Office-Home is available at https://www.hemanthdv.org/officeHomeDataset.html.
-The DomainNet is available at http://ai.bu.edu/M3SDA/.